### v1

# Weight-Entanglement Meets Gradient-Based Neural Architecture Search

Abstract
Weight sharing is a fundamental concept in neural architecture search (NAS), enabling gradient-based methods to explore cell-based architecture spaces significantly faster than traditional blackbox approaches. In parallel, weight entanglement has emerged as a technique for intricate parameter sharing among architectures within macro-level search spaces. Since weight-entanglement poses compatibility challenges for gradient-based NAS methods, these two paradigms have largely developed independently in parallel sub-communities. This paper aims to bridge the gap between these sub-communities by proposing a novel scheme to adapt gradient-based methods for weight-entangled spaces. This enables us to conduct an in-depth comparative assessment and analysis of the performance of gradient-based NAS in weight-entangled search spaces. Our findings reveal that this integration of weight-entanglement and gradient-based NAS brings forth the various benefits of gradient-based methods (enhanced performance, improved supernet training properties and superior any-time performance), while preserving the memory efficiency of weight-entangled spaces. The code for our work is openly accessible here.
1 Introduction
The concept of weight-sharing in Neural Architecture Search (NAS) arose from the need to improve the efficiency of conventional blackbox NAS algorithms, which demand significant computational resources to evaluate individual architectures. Here, weight-sharing (WS) refers to the paradigm by which we represent the search space with a single large supernet, also known as the one-shot model, that subsumes all the candidate architectures in that space. Every edge of this supernet holds all the possible operations that can be assigned to that edge. Importantly, architectures that share a particular operation also share its corresponding operation weights, allowing simultaneous training of an exponential number of subnetworks, unlike the sequential approach of blackbox NAS.
Gradient-based NAS algorithms (or optimizers), such as DARTS (Liu et al., 2019), GDAS (Dong and Yang, 2019) and DrNAS (Chen et al., 2021b), assign an architectural parameter to every choice of operation on a given edge of the supernet. The output feature maps of these edges are thus an aggregation of the outputs of the individual operations on that edge, weighted by their architectural parameters. These architectural parameters are learned using gradient updates by differentiating through the validation loss. Supernet weights and architecture parameters are therefore trained simultaneously in a bi-level fashion. Once this training phase is complete, the final architecture can be identified quickly, e.g., by selecting operations with the highest architectural weights on each edge as depicted in Figure 1(b). However, more sophisticated methods have also been explored (Wang et al., 2021) for this selection.
While gradient-based NAS methods have primarily been studied for cell-based NAS search spaces, a different class of search spaces focuses on macro-level structures (parameterizing kernel size, number of channels, etc.) for which all architectures in the space are subnetworks of the architecture with the largest architectural choices, which is identical to the supernet in this case. These search spaces share weights via the more intricate form of weight-entanglement (WE) between similar operations on the same edge; e.g., the nine weights of a 3 $\times$ 3 convolution are a subset of the 25 weights of a 5 $\times$ 5 convolution. This paradigm reduces the memory requirements of the supernet to the size of the largest architecture in the search space.
In order to efficiently search over such weight-entanglement spaces, two-stage methods have been introduced that first pre-train the supernet, and then perform blackbox search on it to obtain the final architecture. OFA (Cai et al., 2020), SPOS (Guo et al., 2020), AutoFormer (Chen et al., 2021a) and HAT (Wang et al., 2020) are prominent examples of methods that fall into this category. Note that these methods do not employ additional architectural parameters for supernet training or search. They typically train the supernet by randomly sampling subnetworks and training them as depicted in Figure 1(a). The post-hoc blackbox search relies on using the performance of subnetworks sampled from the trained supernet as a proxy for true performance on the unseen test set. To contrast with this two-stage approach, we refer to traditional gradient-based NAS approaches as single stage methods.
While to date, weight-entangled spaces have only been explored with two-stage methods, and cell-based spaces have only been optimized with single-stage approaches, in this paper we bridge the gap between these parallel sub-communities. We do so by addressing the challenges associated with integrating off-the-shelf single-stage NAS methods with weight-entangled search spaces.
After a discussion of related work (Section 2), we make the following main contributions:
  • We propose a generalized scheme to apply single-stage methods to weight-entangled spaces, while maintaining search efficiency and efficacy at larger scales (Section 3, with visualizations in Figure 1(c) and Figure 2). We refer to this method as TangleNAS.
  • We propose a fair playground for the comparative evaluation of single and two-stage methods (Section 4.1) and study the effect of weight-entanglement in conventional cell-based search spaces (i.e., NASBench201 and the DARTS search space) (Section 4.2).
  • We evaluate our proposed generalized scheme for single-stage methods across a diverse set of weight-entangled macro search spaces and tasks, from image classification (Section 4.3.1, Section 4.3.2) to language modeling (Section 4.3.3).
  • We conduct a comprehensive evaluation of the properties of single and two-stage approaches including any-time-performance, memory consumption, robustness to training fraction and effect of fine-tuning (Section 5), demonstrating that our generalized gradient-based NAS method achieves the best of single and two-stage methods: the enhanced performance, improved supernet fine-tuning properties, superior any-time performance of single-stage methods, and the low memory consumption of two-stage methods .
To facilitate reproducibility, our code is openly accessible here.
2 Related Work
Weight-sharing was first introduced in ENAS (Pham et al., 2018), which reduced the computational cost of NAS by 1000$\times$ compared to previous methods. However, since this method used reinforcement learning its performance was quite brittle. Bender et al. (2018) simplified the technique, showing that searching for good architectures is possible by training the supernet directly with stochastic gradient descent. This was followed by DARTS (Liu et al., 2019), which set the cornerstone for efficient and effective gradient-based, single-stage, NAS approaches.
DARTS, however, had prominent failure modes, such as its discretization gap and convergence towards parameter-free operations, as outlined in Robust-DARTS (Zela et al., 2020). Numerous gradient-based one-shot optimization techniques were developed since then (Cai et al., 2019; Nayman et al., 2019; Xu et al., 2019; Dong and Yang, 2019; Hu et al., 2020; Li et al., 2021; Chen et al., 2021b; Zhang et al., 2021). Among these, we highlight DrNAS (Chen et al., 2021b), which we will use in our experiments as a representative of gradient-based NAS methods. DrNAS treats one-shot search as a distribution learning problem, where the parameters of a Dirichlet distribution over architectural parameters are learned to identify promising regions of the search space.
Despite the remarkable performance of single-stage methods, they are not directly applicable to some real-world architectural domains, such as transformers, because of the macro-level structure of these search spaces. DASH (Shen et al., 2022) employs a DARTS-like methodology to optimize CNN topologies (i.e. kernel size, dilation) for a diverse set of tasks, reducing computational complexity by appropriately padding and summing kernels with different sizes and dilations. FBNet-v2 (Wan et al., 2020) makes an attempt along these lines for CNN topologies, but its methodology is not easily extendable to search spaces like transformers with multiple interacting modalities, such as embedding dimension, number of heads, expansion ratio, and depth.
Weight-entanglement, on the other hand, provides a more effective way of weight-sharing exclusive to macro-level architectural spaces. The concept of weight-entanglement was developed in slimmable networks (Yu et al., 2018; Yu and Huang, 2019), OFA (Cai et al., 2020) and BigNAS (Yu et al., 2020) in the context of convolutional networks (see also AtomNAS (Mei et al., 2019)) and later spelled out in AutoFormer (Chen et al., 2021a), where it was applied to the transformer architecture.
Single-path-one-shot (SPOS) methods (Guo et al., 2020) have shown a lot of promise in searching weight-entangled spaces. SPOS trains a supernet by uniformly sampling single paths (one at a time to limit memory consumption) and then training the weights along that path. The supernet training is followed by a black-box search that uses the performance of the models sampled from the trained supernet as a proxy.
OFA used a similar idea to optimize different dimensions of CNN architectures, such as its depth, width, kernel size and resolution. Additionally, it enforced training of larger to smaller sub-networks sequentially to prevent interference between sub-networks. Subsequently, AutoFormer adopted the SPOS method to optimize a weight-entangled space of transformer architectures.
In this work, we demonstrate how single-stage methods can be applied to macro-level search spaces with weight-entanglement, where one can benefit from the time efficiency and remarkable effectiveness of modern differentiable NAS optimizers, while concurrently maintaining memory efficiency of the weight-entangled space. While we chose to adopt DrNAS as the primary approach for our exploration of the weight-entangled space in this study, it is worth highlighting that our methodology is generally applicable to other gradient-based NAS methods.
3 Methodology: Single-Stage NAS with Weight-Superposition
Our primary goal in this work is to effectively apply single-stage NAS to search spaces with macro-level architectural choices. To reduce the memory consumption and preserve computational efficiency, we propose two major modifications to single-stage methods. In our framework, the operators on any given edge inherit the weights of the largest operator on that edge. This reduces the size of the supernet to the size of the largest individual architecture in the search space.
Figure 1: (a) Two-Stage NAS with WE (Algorithm 3): dotted paths show operation choices not sampled at the given step (b) Single-Stage NAS with WS (Algorithm 4): every operation choice is evaluated independently and contributes to the output feature map with corresponding architecture parameters (c) Single-Stage NAS with WE (Algorithm 1): operation choices superimposed with corresponding arhictecture parameters.
Figure 2: Supernet weight matrix (LHS) adapted to gradient-based methods (RHS). All operation choices are zero-padded to match the dimension of the largest operation and superimposed with given architecture parameter. (a) Weight superposition for a single dimension choice (e.g. kernel size) with architecture parameter $\alpha_{i}$ (b) Weight superposition for combination of multiple dimensions (e.g. embedding dimension and expansion ratio) with architecture parameters $\alpha_{i}$, $\beta_{j}$. White areas inside the dashed gray boundaries indicate zero-padding and color values reflect scaling.
Algorithm 1 TangleNAS
1:  Input: $M\leftarrow$ number of cells, $N\leftarrow$ number of operations    $\mathcal{O}\leftarrow[o_{1},o_{2},o_{3},...o_{N}]$    $\mathcal{W}_{\max}\leftarrow\cup_{i-1}^{N}w_{i}$    $\mathcal{A}\leftarrow[\alpha_{1},\alpha_{2},\alpha_{3},...\alpha_{N}]$    $\gamma=$ learning rate of $\mathcal{A}$, $\eta=$ learning rate of $\mathcal{W}_{\max}$   $f$ is a function or distribution s.t. $\sum_{i=1}^{N}f(\alpha_{i})=1$
2:  $Cell_{j}\leftarrow DAG(\mathcal{O}_{j},{\mathcal{W}_{max}}_{j})$ /* defined for j=1...M */
3:  $Supernet\leftarrow\cup_{j}^{M}Cell_{j}\cup\mathcal{A}$
4:  /* example of forward propagation on the cell */
5:  for $j\leftarrow 1$ to $M$ do
6:      /* PAD weight to output dimension of $\mathcal{W}_{\max}$ before summation */
7:      /* Generalized Weighing Scheme */
8:      $\overline{o_{j}(x,\mathcal{W}_{\max})}=o_{j,i}(x,\sum_{i=1}^{N}f(\alpha_{i})\,\mathcal{W}_{\max}[:i])$
9:  end for
10:   /* weights and architecture update */
11:   $\mathcal{A}=\mathcal{A}-\gamma\nabla_{\mathcal{A}}\mathcal{L}_{val}({\mathcal{W}_{\max}}^{*},\mathcal{A})$
12:   $\mathcal{\mathcal{missing}}{W}_{\max}=\mathcal{W}_{\max}-\eta\nabla_{\mathcal{W}_{\max}}\mathcal{L}_{train}(\mathcal{W}_{\max},\mathcal{A})$
13:   /* Architecture Selection */
14:   $selected\_arch\leftarrow$ $\arg\max(\mathcal{A})$
Then, instead of weighting the outputs of operations on a given edge by architectural parameters, as single-stage NAS methods do (see Figure 1(b)), we simply take each operation choice, zero-pad that to match the dimensions of the largest operation and sum these terms, each weighted by the corresponding architectural parameter. Figure 2(a) provides an overview of the idea for a single architectural choice such as the kernel size. This is equivalent to taking the largest operation and re-scaling the weights of each sub-operation by the sum of architectural parameters of operations it is embedded in (see right-most weight matrix in Figure 2(a).
This process results in a superposition of operation weights whose structural property is similar to the largest operation. Consequently, a single forward pass suffices to capture the effect of all operational choices, thus making our approach computationally efficient. Furthermore, to accommodate for operations depending on two or more architectural dimension choices we introduce the combi-superposition outlined in Figure 2(b) and Algorithm 2. This allows us to apply any arbitrary gradient-based NAS method, such as DARTS, GDAS or DrNAS, to macro-level search spaces that leverage weight-entanglement, without additional memory consumption and computational cost during forward propagation. We refer to our memory and compute optimal single-stage architecture search method as TangleNAS and provide an overview of the approach in Algorithm 1. The operation $f$ in Algorithm 1 determines the differentiable optimizer used in the method. A $softmax$ function corresponds to DARTS, while sampling from the Dirichlet distribution corresponds to DrNAS. In all experiments of the upcoming sections we use DrNAS as a basic gradient-based NAS method. We therefore simply refer to this combination as TangleNAS throughout the paper.
4 Experiments
[…]
4.1 Toy search spaces
We begin the evaluation of TangleNAS on two relatively compact toy search spaces that we designed as a contribution to the community to allow faster iterations of algorithm development:
  • Toy cell space: a small version of the DARTS space; architectures are evaluated on the Fashion-MNIST dataset.
  • Toy conv-macro space: a small space inspired by MobileNet, including kernel sizes and the number of channels in each convolution layer; architectures are evaluated on CIFAR-10.
We describe these spaces in Appendix C, including links to code for these open source toy benchmarks. The results of these experiments are summarized in Tables 2 and 2. Across both of these search spaces, TangleNAS outperforms its two-stage counterparts. Also, DrNAS without weight entanglement performs very poorly on the macro level search space.
Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset
Search Type Optimizer Supernet type Test acc Search Time(hrs)
Single-Stage DrNAS WS $91.190\pm 0.049$ 6.277
TangleNAS WE $\mathbf{91.300\pm 0.023}$ 6.222
Two-Stage SPOS+RS WE $90.680\pm 0.253$ 15.611
SPOS+ES WE $90.317\pm 0.223$ 13.244
Optimum - - 91.630 -
Search Type Optimizer Supernet type Test acc Search Time(hrs)
Single-Stage DrNAS WS $10\pm 0.00$ 12.361
TangleNAS WE $\mathbf{82.495\pm.0.461}$ 8.55
Two-Stage SPOS+RS WE $81.253\pm 0.6717$ 21.676
SPOS+ES WE $81.890\pm 0.800$ 26.359
Optimum - - 84.410 -
Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset
Table 2: Evaluation on the toy conv-macro search space on the CIFAR10 dataset. Note that DrNAS with weight sharing converges to a degenerate architecture here.
4.2 Cell-based search spaces
We now start our comparative analysis of single and two-stage approaches by applying them to cell-based spaces that serve as focal points in the single-stage NAS literature to evaluate TangleNAS against DrNAS and SPOS. Here, we use the popular NAS-Bench-201 (NB201) (Dong and Yang, 2020) and DARTS (Liu et al., 2019) search spaces. We refer the reader to Appendix C for details about these spaces and Section E.4 for the experimental setup.
Our contribution on these spaces is two-fold. First, we study the effects of weight-entanglement on cell-based spaces in conjunction with single-stage methods. To this end, we entangle the weights of similar operations with different kernel sizes on both search spaces. For NB201, the weights of the 1x1 and 3x3 convolutions are entangled, and in the DARTS search space, the weights of dilated and separable convolutions with kernel sizes 3x3 and 5x5 are entangled. Second, we study SPOS on the NB201 and DARTS search spaces. To the best of our knowledge, we are the first to study a two-stage method like SPOS in such cell search spaces.
Tables 3 and 4 show the results. For both search spaces, TangleNAS yields the best results, outperforming the single-stage baseline DrNAS with WS, as well as both SPOS variants. TangleNAS also significantly lowers the memory requirements and runtime compared to its weight-sharing counterpart. We note that overall, the SPOS methods are ineffective in these cell search spaces.
Search Type Optimizer Supernet CIFAR10 CIFAR100 ImageNet16-120 Search Time (hrs)
Single-Stage DrNAS WS $\mathbf{94.36\pm 0.00}$ $72.245\pm 0.732$ $\mathbf{46.37\pm 0.00}$ 20.9166
TangleNAS WE $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$ 20.3611
Two-Stage SPOS+RS WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.1460$ 29.51388
SPOS+ES WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.0149$ 26.74721
Table 3: Comparison of test-accuracies of single and two-stage methods on NB201 search space
Search Type Optimizer Supernet CIFAR10 ImageNet Search Time(hrs)
Single-Stage DrNAS WS $2.625\pm 0.075$ 26.29 9.12
TangleNAS WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.691}$ 7.44
Two-Stage SPOS+RS WE $2.965\pm 0.072$ 27.114 18.7444
SPOS+ES WE $3.200\pm 0.065$ 27.320 14.794
Table 4: Comparison of test errors of single and two-stage methods on the DARTS search space
4.3 Macro Search Spaces
Given the promising results of TangleNAS on the toy and cell-based spaces, we now extend our evaluation to the home base of two-stage methods. We study TangleNAS on a vision transformer space (AutoFormer-T) and a convolutional space (MobileNetV3), which have been proposed and explored using two-stage methods by Chen et al. (2021a) and Cai et al. (2020), respectively, as well as a language model transformer search space built around GPT-2 Radford et al. (2019).
4.3.1 AutoFormer
Search Type Optimizer CIFAR10 CIFAR100
Inherit Fine-tune Retrain Params FLOPS Inherit Fine-tune Retrain Params FLOPS
Single-Stage TangleNAS 93.57 $\mathbf{97.702\pm 0.017}$ $\mathbf{97.872\pm 0.054}$ 8.6847M 1.946 76.59 $\mathbf{82.615\pm 0.064}$ $\mathbf{82.668\pm 0.161}$ 8.64874M 1.9388G
Two-Stage SPOS+RS 94.29 $97.605\pm 0.038$ $97.767\pm 0.024$ 8.51161M 1.91G 78.210 $82.407\pm 0.026$ $82.210\pm 0.14242$ 8.47558M 1.9046G
SPOS+ES 94.10 $97.632\pm 0.047$ $97.6425\pm 0.023$ 7.230286M 1.6595 77.97 $82.517\pm 0.140$ $82.5175\pm 0.114$ 8.2447M 1.859G
Table 5: Evaluation on the AutoFormer-T space for CIFAR10 and CIFAR100.
We evaluate TangleNAS on the AutoFormer-T space introduced by Chen et al. (2021a), which is based on vision transformers. The search space consists of the choices of embedding dimensions and number of layers, and for each layer, its MLP expansion ratio and the number of heads it uses to compute attention. More details can be found in Table 19 in the appendix. The embedding dimension choice is held constant throughout the network while the number of heads and the MLP expansion ratio varies for every layer. This amounts to a total search space size of about $10^{13}$ architectures. We train our supernet using the same training hyperparameters as used in AutoFormer.
NAS Method SuperNet-Type ImageNet Datasets Params FLOPS
CIFAR10 CIFAR100 Flowers Pets Cars
SPOS+ES AutoFormer-T 75.474 98.019 86.369 98.066 91.558 91.935 5.893M 1.396G
TangleNAS AutoFormer-T 78.842 98.249 88.290 98.066 92.347 92.396 8.98108M 2.00G
SPOS+ES AutoFormer-S 81.700 99.10 90.459 97.90 94.8529 92.5447 22.9M 5.1G
TangleNAS AutoFormer-S 81.964 99.12 90.459 98.3257 95.07 92.3707 28.806M 6.019G
Table 6: Evaluation on the AutoFormer-T space on downstream tasks. ImageNet accuracies are obtained through inheritance, whereas accuracies for the other datasets are achieved through fine-tuning the imagenet-pretrained model.
Table 5 and Table 6 evaluates TangleNAS against AutoFormer on the AutoFormer-T space. Interestingly, we observe that although AutoFormer sometimes outperforms TangleNAS upon inheritance from the supernet, the TangleNAS architectures are always better upon fine-tuning and much better upon retraining. For ImageNet-1k we obtain an accuracy of $78.842$% by inheriting the weights of the optimal architecture in contrast to $75.474$% by unconstrained evolutionary search on the on AutoFormer-T space (Chen et al., 2021a), i.e. a net improvement of $3.368\%$ (see Table 6).
4.3.2 MobileNetV3
Search Type Optimizer Top-1 acc Params FLOPS
Single-Stage TangleNAS 77.424 7.58M 528.80M
Two-Stage OFA+RS 77.046 6.87M 369.16M
OFA+ES 77.050 7.21M 420.50M
Largest Arch - 77.336 7.66M 566.17M
Table 7: Evaluation on MobileNetV3
Next, we study a convolutional search space based on the MobileNetV3 architecture. The search space is defined in Table 19 in the appendix and contains about $2\times 10^{19}$ architectures. This follows from the search space designed by OFA (Cai et al., 2020), which searches for kernel-size, number of blocks and channel-expansion factor. Unlike OFA, we do not use the progressive shrinking scheme for the supernet training but activate all choices in our supernet at all times.
Table 7 shows that on this OFA search space, TangleNAS still outperforms OFA (based on both unconstrained evolutionary and random search on the pre-trained OFA supernet). Notably, TangleNAS even yields an architecture with higher accuracy than the largest architecture in the space (while OFA yields worse architectures).
4.3.3 Language Modelling Space
Finally, given the growing interest in efficiently training large language models and recent developments in scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), we study our efficient and scalable TangleNAS method on a language-model space.
We create our language model space around a smaller version of Andrej Karpathy’s nanoGPT model [^1] https://github.com/karpathy/nanoGPT. In this transformer search space, we search for the embedding dimension, number of heads, number of layers and MLP ratios as defined in Table 19 (in the appendix). For each of these transformer search dimensions we consider 3 choices. We evaluate our small language model on the TinyStories (Eldan and Li, 2023) dataset. We improve over the SPOS+ES and SPOS+RS baselines as presented in Table 8. This improvement is statistically very significant with a two-tailed p-value of $0.0064$ for ours v/s SPOS+ES and p-value of $0.0127$ for ours v/s SPOS+RS.
[…]
Single-Stage TangleNAS $\mathbf{1.412\pm 0.011}$ 4.104 87.01M 93.88s
Two-Stage SPOS+RS $1.433\pm 0.005$ 4.191 77.57M 85.169s
SPOS+ES $1.444\pm 0.013$ 4.238 78.75M 87.10s
Table 8: Comparison of single and two stage methods on language model search space. We report the test loss and perplexity on the tinystories dataset
5 Results and Discussion
For a more in-depth evaluation, we now compare different properties of single-stage and two-stage methods, starting from their any-time performance, impact of the portion of train-val split on their performance and centered-kernel alignment of the supernet feature maps. Moreover, we also study the effect of pretraining, fine-tuning and retraining on the AutoFormer space for CIFAR10 and CIFAR100 and downstream performance of imagenet pre-trained best model on several classification datasets. Finally, we discuss the insights derived from our single-stage approach in designing architectures on real world tasks.
Space and time complexity
In practice, we observe that vanilla gradient-based NAS methods are memory and compute expensive in comparison to both two-stage methods and our TangleNAS approach with weight-superposition (described in Figure 2). In particular on the NB201 and DARTS search spaces we observe a 25.276% and 35.545% reduction in memory requirements for TangleNAS over DrNAS with WS. This issue only exacerbates for weight-entangled spaces like AutoFormer and MobileNetV3, making the application of vanilla gradient-based methods practically infeasible. This is because vanilla gradient-based methods incur an $\mathcal{O}(n)$ overhead over TangleNAS in terms of space and time complexity, where n is the number of operation choices.
Any-Time performance
An essential characteristic of any neural architecture search (NAS) method is the attainment of strong any-time performance. End users of NAS techniques frequently prioritize the expedited identification of competitive architectures within limited time constraints. This becomes particularly crucial in light of the escalating computational expenses associated with training sophisticated and large neural networks, such as Transformers. Strong any-time performance is a pivotal criterion for the effective applicability of NAS in such resource-intensive domains. We thus compare the any-time performance of TangleNAS. Figure 3 demonstrates that TangleNAS (DrNAS+WE) is faster than its baseline method DrNAS+WS. Similarly, Figure 4 shows that TangleNAS has better anytime performance than AutoFormer.
(a) CIFAR100
(b) ImageNet16-120
Figure 3: Test accuracy evolution over epochs for NB201
(a) CIFAR10
(b) CIFAR100
Figure 4: Any Time performance curves of AutoFormer vs. Ours
Figure 5: MLP ratio trajectory for LM. Number of layers range from 1-12 and mlp ratio choice can be 2,3 or 4.
Effect of fraction of training data
One-shot NAS commonly employs a 50%-50% train-valid split for cell-based spaces and 80%-20% for weight-entangled spaces. To avoid potential confounding factors, we also study our method across different data splits for each search space. The results reported in Appendix A are largely similar across splits. In particular, we observe that single-stage methods are quite robust and performant across training fractions and search spaces, in comparison to two-stage methods.
Architecture design insights
In transformer spaces reducing the mlp-ratio in the initial layers has a relatively low impact on performance (Figure 5) and can often work competitively or outperform handcrafted architectures. This observation is consistent across ViT and Language Model spaces. Conversely, number of heads and embedding dimension have a significant impact on the quantitative metric of choice. Pruning a few of the final layers also has a relatively low impact on performance. In the MobileNetV3 space, we find a strong preference for larger number of channels and larger network depth. In contrast, we discover that scaling laws for transformers may not necessary apply in convolutional spaces especially for kernel sizes. We observe that earlier layers favor 5x5 kernel sizes while later ones prefer 7x7 kernel sizes (3 being the smallest and 7 the largest).
Effect of pretraining, fine-tuning and retraining
We study the impact of inheriting, fine-tuning, and retraining on the CIFAR10 and CIFAR100 datasets on the AutoFormer space. We observe that retraining almost always outperforms fine-tuning and inheriting. This brings to question the rank correlation between the inherited accuracy and retraining accuracy of two-stage methods and the training interference (Xu et al., 2022) in two-stage methods. Strong rank-correlation is desirable especially in two-stage methods as these methods rely on the performance proxy upon inheritance for the black-box search. Furthermore, we observe that while the SPOS+RS and SPOS+ES methods are performant upon inheritance, TangleNAS outperforms these models after fine-tuning and training from scratch. Improved supernet training properties for single-stage methods is further supported by Centered Kernel Alignment (CKA) analysis (Kornblith et al., 2019) (Table 25).
ImageNet-1k pre-trained architecture on downstream tasks
Finally, we study the impact on fine-tuning the best model obtained from the search on downstream datasets. We follow the fine-tuning pipeline proposed in AutoFormer and fine-tune on different fine and coarse-grained datasets. We observe from Table 6 that the architecture discovered by TangleNAS on ImageNet is much more performant in fine-tuning to various datasets (CIFAR-10, CIFAR-100, Flowers, Pets and Cars) than the architecture discovered by SPOS.
6 Conclusion
In this paper, we investigate two of the most widely used NAS methods: single-stage and two-stage, each traditionally applied to different types of search spaces. We extend the applicability of single-stage NAS methods to weight-entangled spaces, a domain primarily explored by two-stage methods. By doing so, we aim to bridge the gap between these two distinct sub-fields. We empirically evaluate our proposed single-stage method, TangleNAS, on a diverse set of weight-entangled search spaces and tasks, showcasing its ability to outperform conventional two-stage NAS methods while enhancing search efficiency. Furthermore, we highlight the advantages of introducing weight-entanglement into traditional cell-based spaces, whenever feasible. Doing so not only improves the speed of single-stage methods but also reduces memory consumption. The modular nature of TangleNAS enables it to accommodate any differentiable optimizer and hence take advantage of the continually-improving set of methods for gradient-based NAS. Our results on various macro-level search spaces are already very encouraging and we therefore hope that this line of work will enable the NAS community to further improve modern architectures like Transformers.
6.0.1 Acknowledgments
This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the German Federal Ministry of Education and Research (BMBF, grant RenormalizedFlows 01IS19077C); the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Robert Bosch GmbH is acknowledged for financial support. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
Appendix A Training across data fractions
Search Type Optimizer Train portion CIFAR10 CIFAR100
Single-Stage TangleNAS 50% $97.715\pm 0.0877$ $82.5375\pm 0.11764$
80% $\mathbf{97.872\pm 0.05369}$ $\mathbf{82.6675\pm 0.16073}$
Two-Stage SPOS+RS 50% $97.6799\pm 0.0262$ $82.5372\pm 0.27954$
80% $97.7674\pm 0.0239278$ $82.210\pm 0.14242$
SPOS+RE 50% $97.77\pm 0.03752$ $82.354\pm 0.11974$
80% $97.6425\pm 0.02286$ $82.5175\pm 0.1135$
Table 9: Evaluation on the AutoFormer-T space for CIFAR10 and CIFAR100
Search Type Optimizer Train portion Supernet Accuracy (%)
CIFAR10 CIFAR100 ImageNet16-120
Single-Stage DrNAS 50% WS $\mathbf{94.36\pm 0.00}$ $72.245\pm 0.7315$ $\mathbf{46.37\pm 0.00}$
80% $\mathbf{94.36\pm 0.00}$ $71.1525\pm 0.697$ $\mathbf{46.37\pm 0.00}$
TangleNAS 50% WE $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$
80% $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$
Two-Stage SPOS+RS 50% WE $89.107\pm 0.88392$ $56.865\pm 2.5968$ $31.665\pm 1.1460$
80% $87.778\pm 2.4462$ $53.68\pm 4.1736$ $30.5449\pm 3.6425$
SPOS+ES 50% WE $87.133\pm 2.6050$ $56.463\pm 2.3417$ $29.7849\pm 3.0149$
80% $89.095\pm 0.8250$ $56.363\pm 4.7244$ $30.935\pm 3.54648$
Table 10: Comparison of test accuracies of single and two stage methods with WS and WE on NB201 search space
Table 11: Evaluation on toy cell-based search space on fashion-mnist dataset
Search Type Optimizer Train portion Supernet type Test acc
Single-Stage DrNAS 50% WS $91.19\pm 0.0490$
80% $91.125\pm 0.0334$
TangleNAS 50% WE $\mathbf{91.3\pm 0.023}$
80% $91.065\pm 0.1633$
Two-Stage SPOS+RS 50% WE $90.68\pm 0.25345$
80% $90.687\pm 0.10969$
SPOS+ES 50% WE $90.3175\pm 0.2233$
80% $90.595\pm 0.2188$
Optimum - - - 91.63
Search Type Optimizer Train portion Supernet type Test acc
Single-Stage DrNAS 50% WS 10%
80% 10%
TangleNAS 50% WE $\mathbf{83.020\pm 0.000}$
80% 82.495±.0.46074plus-or-minus82.495.0.4607482.495$\pm$.0.4607482.495 ± .0.46074
Two-Stage SPOS+RS 50% WE $81.253\pm 0.67174$
80% $81.345\pm 0.38336$
SPOS+ES 50% WE $81.890\pm 0.8002$
80% $82.322\pm 0.60356$
Optimum - - - 84.41
Table 11: Evaluation on toy cell-based search space on fashion-mnist dataset
Table 12: Evaluation on toy conv-macro search space on CIFAR10 dataset
Search Type Optimizer Train portion Supernet CIFAR10 ImageNet
Single-Stage TangleNAS 50% WE $\mathbf{2.556\pm 0.03406}$ $\mathbf{25.69}$
80% $2.67\pm 0.0756$ 25.742
DrNAS 50% WS $2.625\pm 0.0750$ 26.29
80% $2.580\pm 0.02798$ $\mathbf{25.67}$
Two-Stage SPOS+RS 50% WE $2.965\pm 0.07193$ 27.114
80% $2.965\pm 0.07193$ 27.114
SPOS+RE 50% WE $3.20000\pm 0.06557$ 27.424
80% $3.00249\pm 0.037052$ 26.76
Table 13: Comparison of test errors of single and two stage methods with WS and WE on DARTS search space
Appendix B Search and Training Times
B.1 NB201 and DARTS
Search Type Optimizer Supernet Search Time (hrs)
Single-Stage DrNAS WS 20.9166
TangleNAS WE 20.3611
Two-Stage SPOS+RS WE 29.51388
SPOS+ES WE 26.74721
Table 14: Comparison of search times NB201
Search Type Optimizer Supernet GPU MemSearch Time(hrs)
Single-Stage DrNAS WS 81.08% 29.4444
TangleNAS WE 52.26% 25.472
Two-Stage SPOS+RS WE 25.66% 18.7444
SPOS+ES WE 25.66% 14.794
Table 15: Comparison of test errors of single and two-stage methods on the DARTS search space
Appendix C Search Space details
Toy cell space
This particular search space takes its inspiration from the prominently used Differentiable Architecture Search (DARTS) space, and is composed of diminutive triangular cells, with each edge offering four choices of operations: (a) Separable 3x3 Convolution, (b) Separable 5x5 Convolution, (c) Dilated 3x3 Convolution, and (d) Dilated 5x5 Convolution. The macro-architecture of the model comprises of three cells of types reduction, normal, and reduction again stacked together. Notably, we entangle the 3x3 and 5x5 kernel weights for each operation type, i.e., separable convolutions and dilated convolutions. We evaluate these search spaces and their architectures on the Fashion-MNIST dataset by creating a small benchmark which we release here.
Toy conv-macro space
This toy space draws its inspiration from MobileNet-like spaces where we search for the number of channels and the kernel size of convolutional layers in a network (also called macro search space) for four convolutional layers. Every convolutional layer has a choice of three kernel sizes and number of channels. See Table 16 for more details. We evaluate this search space on the CIFAR10 dataset by creating a small benchmark which we release here.
Architectural Parameter Choices
Kernel sizes (all layers) 3, 5, 7
Number of channels (layer 1) 8, 16, 32
Number of channels (layer 2) 16, 32, 64
Number of channels (layer 3) 32, 64, 128
Number of channels (layer 4) 64, 128, 256
Table 16: Toy Convolutional-Macro Search Space
NASBench201 Space
The NASBench201 [Dong and Yang, 2020] has a single cell type which is stacked 5 times. Each cell has 4 nodes and each node are connected by operations from amongst skip connection, conv 3x3, conv 1x1, max-pool or average-pool. Opportunities for weight entanglement in this space is limited and we entangle the conv 3x3 and conv 1x1 kernel on every edge for every cell, thus obtaining parameter savings over traditional weight sharing.
DARTS Search Space
The DARTS [Liu et al., 2019] search space has a cell with 13 possible edges stacked 8 times. The nodes of the cell are connected by operations amongst :none, max_pool_3x3, avg_pool_3x3, skip_connect, sep_conv_5x5, sep_conv_3x3, dil_conv_5x5, dil_conv_3x3.
AutoFormer Space and Language Model Space
We present the details of the AutoFormer Space and the Language Model space is Table 19 and Table 19 respectively.
Table 17: Choices for AutoFormer-T Search Space
Table 18: Choices for Language Model Space
Architectural Parameter Choices
Embedding dimension 192, 216, 240
Number of layers 12, 13, 14
MLP ratio (per layer) 3.5, 4
Number of heads (per layer) 3, 4
Architectural Parameter Choices
Embedding Dimension 384, 576, 768
Number of Heads 6, 8, 12
MLP Expansion Ratio 2, 3, 4
Number of Layers 5,6,7
Architectural Parameter Choices
Kernel sizes (all layers) 3, 5, 7
Channel expansion (all layers) 3,4,6
Number of blocks 2,3,4
Table 17: Choices for AutoFormer-T Search Space
Table 18: Choices for Language Model Space
Table 19: MobileNet Search Space
Appendix D Details on Method
Traditional cell-based search spaces primarily consider independent operations (e.g., convolution or skip). One-shot differentiable optimizers thus have their mixture operations tailored to these search spaces which are not general enough to be applied to macro-level architectural parameters. Consider, e.g., the task of searching for the embedding dimension and the expansion ratio for a transformer. Here, a single operation, i.e., a linear expansion layer, has two different architectural parameters - one corresponding to the choice of embedding dimension and the other to the expansion ratio. To adapt single-stage methods to these combined operation choices in the search space, we propose the combi-superposition operation.
The combi-superposition operation simply takes the cross product of architectural parameters for the embedding dimension and expansion ratio and assigns its elements to every combination of these dimensions, allowing us to optimize jointly in this space without the need for a separate forward pass for each combination. Every combination maps to a unique sub-matrix of the operator weight matrix, indexed using both the embedding dimension and the expansion ratio. To address shape mismatches of the different operation weights during forward passes, every sub-matrix is zero-padded to match the shape of the largest matrix. See Algorithm 2 for more details, and Figure 2(b) for an overview of the idea.
Algorithm 2 Combi-Superposition Operation
  $embed\_dim\leftarrow[e_{1},e_{2},e_{3},\ldots,e_{n}]$Choices for embedding dimension}
  $expansion\_ratio\leftarrow[r_{1},r_{2},r_{3},\ldots,r_{m}]$Choices for expansion ratio}
  $\alpha\leftarrow[\alpha_{1},\alpha_{2},\alpha_{3},\ldots,\alpha_{n}]$Architecture parameters for embedding dimension}
  $\beta\leftarrow[\beta_{1},\beta_{2},\beta_{3},\ldots,\beta_{m}]$Architecture parameters for expansion ratio}
  $X\leftarrow input\_feature$
  $W,b\leftarrow fc\_layer\_weight,fc\_layer\_bias$
  $W_{mix}\leftarrow\textbf{0}$
  $b_{mix}\leftarrow\textbf{0}$
  $\alpha,\beta\leftarrow\text{normalize}(\alpha),\text{normalize}(\beta)$Normalize architecture parameters (e.g., using softmax)}
  $\alpha\beta\leftarrow\alpha\times\beta$Cross product of $\alpha$ and $\beta$ for dimension $n\times m$}
  for $i\leftarrow 1$ to $n\times m$ do
     $W\_i=W[:(embed\_dim[i]\times expansion\_ratio[i]),:embed\_dim[i]]$
     $b\_i=b[:embed\_dim[i]\times expansion\_ratio[i]]$
     $W_{mix}=W_{mix}+\alpha\beta[i]\times\text{PAD}(W_{i})$
     $b_{mix}=b_{mix}+\alpha\beta[i]\times\text{PAD}(b_{i})$
  end for
  $Y\leftarrow X\cdot W_{mix}+b_{mix}$Compute the output of the FC layer with a mixture of weights and bias}
  return  $Y$
For completeness and for comparison with TangleNAS (Algorithm 1), we present Algorithm 3 which describes a generic two-stage method on a macro search space with WE. Also, vanilla single-stage methods on cell-based WS spaces follow Algorithm 4.
Algorithm 3 Weight Entanglement (Two-Stage)
1:  Input: $M\leftarrow$ number of cells, $N\leftarrow$ number of operations    $\mathcal{O}\leftarrow[o_{1},o_{2},o_{3},...o_{N}]$   $\mathcal{W}_{\max}\leftarrow\cup_{i-1}^{N}w_{i}$    $\eta=$ learning rate of $\mathcal{W}_{\max_{\mathcal{O}}}$
2:  $Cell_{j}\leftarrow DAG(\mathcal{O_{j},\mathcal{W}_{\max_{j}}})$ /* defined for j=1...M */
3:  $Supernet\leftarrow\cup_{i}^{M}Cell_{i}$
4:  /* example of forward propagation on the cell */
5:  for $j\leftarrow 1$ to $M$ do
6:      $i\sim\mathcal{U}(1,N)$
7:      /* Slice weight matrix corresponding to operation */
8:      $o_{j}(x,\mathcal{W}_{max})=o_{(j,i)}(x,\mathcal{W}_{max}[:i])$
9:  end for
10:   /* weights update */
11:   $\mathcal{W}_{\max}[:i]=\mathcal{W}_{\max}[:i]-\eta\nabla_{\mathcal{W}_{\max}}[:i]\mathcal{L}_{train}(\mathcal{W}_{\max})$
12:   /* Search */ 13:  $Supernet^{*}\leftarrow$ pre-trained supernet 14:  $selected\_arch\leftarrow$ Evolutionary-Search($Supernet^{*}$)
Algorithm 4 Weight Sharing (Single-Stage)
1:  Input: $M\leftarrow$ number of cells, $N\leftarrow$ number of operations    $\mathcal{O}\leftarrow[o_{1},o_{2},o_{3},...o_{N}]$    $\mathcal{W_{O}}\leftarrow[w_{1},w_{2},w_{3},....w_{N}]$    $\mathcal{A}\leftarrow[\alpha_{1},\alpha_{2},\alpha_{3},...\alpha_{N}]$    $\gamma=$ learning rate of $\mathcal{A}$   $\eta=$ learning rate of $\mathcal{W_{O}}$    $f$ is a function or distribution s.t. $\sum_{i=1}^{N}f(\alpha_{i})=1$
2:  $Cell_{i}\leftarrow DAG(\mathcal{O_{j},\mathcal{W_{O_{j}}}})$ /* defined for i=1...M */
3:  $Supernet\leftarrow\cup_{i}^{M}Cell_{i}\cup\mathcal{A}$
4:  /* example of forward propagation on the cell */
5:  for $j\leftarrow 1$ to $M$ do
6:      /* Compute mixture operation as weighted sum of output of operations*/
7:      $\overline{o_{j}(x,\mathcal{W_{O}})}=\sum_{i=1}^{N}f(\alpha_{i})\,o_{(j,i)}(x,w_{(j,i)})$
8:  end for
9:   /* weights and architecture update */
10:   $\mathcal{A}=\mathcal{A}-\gamma\nabla_{\mathcal{A}}\mathcal{L}_{val}({\mathcal{W_{O}}}^{*},\mathcal{A})$
11:   $\mathcal{W_{O}}=\mathcal{W_{O}}-\eta\nabla_{\mathcal{W_{O}}}\mathcal{L}_{train}(\mathcal{W_{O}},\mathcal{A})$
12:   /* Search */
13:   $selected\_arch\leftarrow$ $\arg\max(\mathcal{A})$
Appendix E Experimental Setup
E.1 Toy Search Spaces
Search Space Cell Space Conv Macro
Epochs 100 100
Learning rate (LR) 0.1 3e-4
Min. LR 0.001 0.0001
Optimizer SGD Adam
Architecture LR 0.0003 0.0003
Batch Size 64 64
Momentum 0.9 0.9
Nesterov True False
Weight Decay 0.0005 0.0005
Arch Weight Decay 0.001 0.001
Regularization Type L2 L2
Regularization Scale 0.001 0.001
Table 20: Configurations used in the DrNAS experiments on Toy Spaces.
Search Space Cell Space Conv Macro
Epochs 250 250
Learning rate (LR) 0.1 3e-4
Min. LR 0.001 0.0001
Optimizer SGD Adam
Batch Size 64 64
Momentum 0.9 0.9
Nesterov True False
Weight Decay 0.0005 0.0005
Table 21: Configurations used in the SPOS experiments on Toy Spaces.
E.2 Language Model
We use the AdamW optimizer in all experiments pertaining to language modelling. Other hyperparameter choices are as in Table 22.
Search Space Small-LM
Learning rate (LR) 5e-4
Min LR 5e-5
Beta2 0.99
Warmup Iters 100
Max Iters 6000
Lr decay iters 6000
Batch size 12
Weight decay 1e-1
Table 22: Configurations used in DrNAS on the Language Model Spaces
E.3 AutoFormer and OFA
We use a 50%-50% train and validation split for the CIFAR10 and CIFAR100 dataset for the cell-based spaces and a 80%-20% for the weight-entangled spaces We use the official source code of AutoFormer available at code for all the AutoFormer experiments on CIFAR10 and CIFAR100. We closely follow the AutoFormer training pipeline and search space design. AutoFormer searched on three transformer sizes Autoformer-T (tiny), AutoFormer-S (small) and Autoformer-B (base). We currently restrict ourselves to Autoformer-T. For baselines like OFA and AutoFormer we follow their respective recipes to obtain the train-valid split for ImageNet-1k. Our models were trained on 2xA100s with the same effective batch-size as AutoFormer. For MobileNetV3 from once-for-all we use the same training hyperparameters as the baseline as here (in addition to architectural parameters same as Table 23) .
E.3.1 AutoFormer Fine-tuning
C10 and C100 pretrained supernet
We fine-tune the C10 and C100 selected networks (after inheriting them from the supernet) for 500 epochs. We set the learning rate to 1e-3, the warmup epochs to 5 , the warmup learning rate to 1e-6 and the min learning rate to 1e-5. All other hyperparameters are set same as Section E.3.
ImageNet pretrained supernet
We follow the DeiT [Touvron et al., 2021] finetuning pipeline as used in AutoFormer. Mainly we set the epochs to 1000, the warmup-epochs to 5 , the scheduler to cosine, the mixup to 0.8, the smoothing to 0.1, the weight-decay to 1e-4, the batch-size 64, the optimizer to sgd, the learning rate to 0.01 and the warmup-lr to 0.0001 for all datasets.
E.4 NB201 and DARTS
For single-stage optimizers, the supernet was trained with four different seeds. The supernet with the best validation performance from these four was discretized to obtain the final model, which is then trained from scratch four times to obtain the results shown in the table. For two-stage methods, yet again, we train the supernet four times, and perform Random Search (RS) and Evolutionary Search (ES) on each of them. The best model obtained across all four supernets for both methods are then trained from scratch with four seeds to compute the final results. For each DrNAS we follow the same recipe as suggested by the authors across all search spaces. To accommodate for multiple training recipes, we write a configurable training pipeline. The configurations for DrNAS, and SPOS are shown in Table 23 and Table 24, respectively.
Search Space DARTS NB201
Epochs 50 100
Learning rate (LR) 0.1 0.025
Min. LR 0.0 0.001
Architecture LR 0.0006 0.0003
Batch Size 64 64
Momentum 0.9 0.9
Nesterov True False
Weight Decay 0.0003 0.0003
Arch Weight Decay 0.001 0.001
Partial Connection Factor 6 -
Regularization Type L2 L2
Regularization Scale 0.001 0.001
Table 23: Configurations used in the DrNAS experiments.
Search Space DARTS NB201
Epochs 250 250
Learning rate (LR) 0.025 0.025
Min. LR 0.001 0.001
Architecture LR 0.0003 0.0003
Batch Size 256 64
Momentum 0.9 0.9
Nesterov True True
Weight Decay 0.0005 0.0005
Arch Weight Decay 0.001 0.001
Table 24: Configurations used in the SPOS experiments.
Appendix F Optimal architectures derived
F.1 DARTS
Figure 6: DRNAS with WE normal cell
Figure 7: DRNAS with WE reduction cell
F.2 Small LM
num_layers: 7, embed_dim: 768, num_heads: [12, 12, 12, 12, 12, 12, 12], mlp_ratio: [3, 4, 4, 4, 4, 4, 4]
F.3 AutoFormer
F.3.1 CIFAR10
50%50% split
mlp_ratio:$[4,4,4,4,4,4,4,4,4,3.5,4,4,3.5,4]$, num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4]44444444444444[4,4,4,4,4,4,4,4,4,4,4,4,4,4][ 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 ], num_layers: 14 embed_dim: 216
80%-20% split
mlp_ratio:$[4,4,4,4,4,4,4,4,3.5,4,4,4,3.5,3.5]$, num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4]44444444444444[4,4,4,4,4,4,4,4,4,4,4,4,4,4][ 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 ], depth: 14, embed_dim: 240
F.3.2 CIFAR100
50%-50% split
mlp_ratio: [4,4,4,4,4,4,4,4,3.5,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4], depth: 14, embed_dim: 216
80%-20% split
mlp_ratio:[3.5,4,4,4,4,4,4,4,4,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4, depth: 14, embed_dim: 240
F.3.3 ImageNet1-k
mlp_ratio:[4,4,4,4,4,4,4,4,4,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4, depth: 14, embed_dim: 240
F.4 MobileNetV3
Kernel_sizes:[7,5,5,7,5,5,7,7,5,7,7,7,5,7,7,5,5,7,7,5],Channel_expansion_factor:[6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6], Depths : [4, 4, 4, 4, 4]
F.5 Toy Spaces
F.5.1 Toy cell (our best architecture)
50%-50% split
Genotype(normal=[(’dil_conv_3x3’, 0), (’dil_conv_3x3’, 0), (’sep_conv_3x3’, 1)], normal_concat=range(1, 3), reduce=[(’sep_conv_3x3’, 0), (’sep_conv_3x3’, 0), (’dil_conv_3x3’, 1)], reduce_concat=range(1, 3))
80%-20% split
Genotype(normal=[(’dil_conv_3x3’, 0), (’dil_conv_5x5’, 0), (’sep_conv_3x3’, 1)], normal_concat=range(1, 3), reduce=[(’sep_conv_3x3’, 0), (’sep_conv_5x5’, 0), (’dil_conv_3x3’, 1)], reduce_concat=range(1, 3))
F.6 Toy conv-macro (our best architecture)
50%-50% split:
Channels = [32, 64, 128, 64], Kernel Sizes = [5, 5, 7, 7].
Train-Val fraction 80%-20%:
Channels = [32, 64, 128, 64], Kernel Sizes = [5, 5, 7, 7].
Appendix G Architecture Representation Analysis
CKA
Centered Kernel Alignment (CKA) [Kornblith et al., 2019] is a metric, based on the Hilbert- Schmidt Independence Criterion (HSIC). This metric is primarily designed to model similarity between representations in neural networks. In this section we study the CKA between structurally identical layers in the inherited, fine-tuned and retrained networks in the AutoFormer-T space. Precisely, the goal here is to study how similar the inherited and fine-tuned representations are to the actual model trained from scratch. Table 25 Presents the average CKA values for a fixed subsample of the CIFAR10 and CIFAR100 dataset. We observe that the representations learned by the single-stage supernet are more similar to the architecture finetuned or trained from scratch. This further emphasizes the potential issues with using the inherited weights from the supernet as proxy for search for two-stage methods as highlighted in [Xu et al., 2022].
Model Inherit v/s Retrain Fine-Tune v/s Retrain
CIFAR10 CIFAR100 CIFAR10 CIFAR100
TangleNAS 0.4630 0.5853 0.57125 0.65275
SPOS+ES 0.45812 0.57932 0.569374 0.6309
SPOS+RS 0.44124 0.583 0.5797 0.638948
Table 25: CKA correlation between layers
Architecture Search-Type Loss Perplexity Params Inference Time
GPT-2 Manual 3.0772 21.69 123.59M 113.3s
TangleNAS Automated 2.9038 18.243 116.51M 102.5s
  1. 1.
    Best Practices for Releasing Code
    For all experiments you report:
    1. (a)
      Did you release code for the training pipeline used to evaluate the final architectures? Yes
    2. (b)
      Did you release code for the search space? Yes
    3. (c)
      Did you release the hyperparameters used for the final evaluation pipeline, as well as random seeds? Yes
    4. (d)
      Did you release code for your NAS method? Yes
    5. (e)
      Did you release hyperparameters for your NAS method, as well as random seeds? Yes
  2. 2.
    Best practices for comparing NAS methods
    1. (a)
      For all NAS methods you compare, did you use exactly the same NAS benchmark, including the same dataset (with the same training-test split), search space and code for training the architectures and hyperparameters for that code? Yes
    2. (b)
      Did you control for confounding factors (different hardware, versions of DL libraries, different runtimes for the different methods)? Yes
    3. (c)
      Did you run ablation studies? Yes
    4. (d)
      Did you use the same evaluation protocol for the methods being compared? Yes
    5. (e)
      Did you compare performance over time? Yes
    6. (f)
      Did you compare to random search? Yes. Black box random search is very expensive in our case, hence we use the supernet from single-path-one-shot(SPOS) for random search.
    7. (g)
      Did you perform multiple runs of your experiments and report seeds? Yes
    8. (h)
      Did you use tabular or surrogate benchmarks for in-depth evaluations? Yes
  3. 3.
    Best practices for reporting important details
    1. (a)
      Did you report how you tuned hyperparameters, and what time and resources this required? N/A. We did not tune any hyperparameters.
    2. (b)
      Did you report the time for the entire end-to-end NAS method (rather than, e.g., only for the search phase)? Yes
    3. (c)
      Did you report all the details of your experimental setup? Yes
Eer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@
 Abstract
-Weight sharing is a fundamental concept in neural architecture search (NAS), enabling gradient-based methods to explore cell-based architecture spaces significantly faster than traditional blackbox approaches.
-In parallel, weight entanglement has emerged as a technique for intricate parameter sharing among architectures within macro-level search spaces.
-Since weight-entanglement poses compatibility challenges for gradient-based NAS methods, these two paradigms have largely developed independently in parallel sub-communities.
+Weight sharing is a fundamental concept in neural architecture search (NAS), enabling gradient-based methods to explore cell-based architectural spaces significantly faster than traditional black-box approaches.
+In parallel, weight-entanglement has emerged as a technique for more intricate parameter sharing amongst macro-architectural spaces.
+Since weight-entanglement is not directly compatible with gradient-based NAS methods, these two paradigms have largely developed independently in parallel sub-communities.
 This paper aims to bridge the gap between these sub-communities by proposing a novel scheme to adapt gradient-based methods for weight-entangled spaces.
 This enables us to conduct an in-depth comparative assessment and analysis of the performance of gradient-based NAS in weight-entangled search spaces.
-Our findings reveal that this integration of weight-entanglement and gradient-based NAS brings forth the various benefits of gradient-based methods (enhanced performance, improved supernet training properties and superior any-time performance), while preserving the memory efficiency of weight-entangled spaces.
+Our findings reveal that this integration of weight-entanglement and gradient-based NAS brings forth the various benefits of gradient-based methods, while preserving the memory efficiency of weight-entangled spaces.
 The code for our work is openly accessible here.
 1 Introduction
-The concept of weight-sharing in Neural Architecture Search (NAS) arose from the need to improve the efficiency of conventional blackbox NAS algorithms, which demand significant computational resources to evaluate individual architectures.
+The concept of weight-sharing in Neural Architecture Search (NAS) was motivated by the need to improve efficiency over that of conventional black-box NAS algorithms, which demand significant computational resources to evaluate individual architectures.
 Here, weight-sharing (WS) refers to the paradigm by which we represent the search space with a single large supernet, also known as the one-shot model, that subsumes all the candidate architectures in that space.
 Every edge of this supernet holds all the possible operations that can be assigned to that edge.
-Importantly, architectures that share a particular operation also share its corresponding operation weights, allowing simultaneous training of an exponential number of subnetworks, unlike the sequential approach of blackbox NAS.
-Gradient-based NAS algorithms (or optimizers), such as DARTS (Liu et al., 2019), GDAS (Dong and Yang, 2019) and DrNAS (Chen et al., 2021b), assign an architectural parameter to every choice of operation on a given edge of the supernet.
+Importantly, architectures that share a particular operation also share its corresponding operation weights, allowing for efficient simultaneous partial training of an exponential number of subnetworks with each gradient update.
+Gradient-based NAS algorithms (or optimizers), such as DARTS (liu-iclr19a), GDAS (dong-cvpr19a) and DrNAS (chen-iclr21a), assign an architectural parameter to every choice of operation on a given edge of the supernet.
 The output feature maps of these edges are thus an aggregation of the outputs of the individual operations on that edge, weighted by their architectural parameters.
@@
 Supernet weights and architecture parameters are therefore trained simultaneously in a bi-level fashion.
-Once this training phase is complete, the final architecture can be identified quickly, e.g., by selecting operations with the highest architectural weights on each edge as depicted in Figure 1(b).
-However, more sophisticated methods have also been explored (Wang et al., 2021) for this selection.
-While gradient-based NAS methods have primarily been studied for cell-based NAS search spaces, a different class of search spaces focuses on macro-level structures (parameterizing kernel size, number of channels, etc.) for which all architectures in the space are subnetworks of the architecture with the largest architectural choices, which is identical to the supernet in this case.
-These search spaces share weights via the more intricate form of weight-entanglement (WE) between similar operations on the same edge; e.g., the nine weights of a 3 $\times$ 3 convolution are a subset of the 25 weights of a 5 $\times$ 5 convolution.
-This paradigm reduces the memory requirements of the supernet to the size of the largest architecture in the search space.
-In order to efficiently search over such weight-entanglement spaces, two-stage methods have been introduced that first pre-train the supernet, and then perform blackbox search on it to obtain the final architecture.
-OFA (Cai et al., 2020), SPOS (Guo et al., 2020), AutoFormer (Chen et al., 2021a) and HAT (Wang et al., 2020) are prominent examples of methods that fall into this category.
+Once this training phase is complete, the final architecture can be obtained quickly, e.g., by selecting operations with the highest architectural weights on each edge as depicted in Figure 1(b).
+However, more sophisticated methods have also been explored (wang-iclr21a) for this selection.
+While gradient-based NAS methods have primarily been studied for cell-based search spaces, a different class of search spaces focuses on macro-level structural decisions, such as the number of channels in a layer of the supernet, or the number of layers which are stacked to form the supernet.
+In these spaces, all architectures are subnetworks of the architecture with the largest architectural choices, identical to the supernet in this case.
+These search spaces share weights more intricately via weight-entanglement (WE) between similar operations on the same edge.
+An example of this for convolutional layers is that the 9 weights of a 3 $\times$ 3 convolution are a subset of the 25 weights of a 5 $\times$ 5 convolution.
+This paradigm reduces the memory footprint of the supernet to the size of the largest architecture in the space, unlike WS search spaces, where it increases linearly with the number of operation choices.
+In order to efficiently search over such weight-entanglement spaces, two-stage methods first pre-train the supernet and then perform black-box search on it to obtain the final architecture.
+OFA (cai-iclr20a), SPOS (guo-eccv20a), AutoFormer (chen-iccv21b) and HAT (wang-acl20) are prominent examples of methods that fall into this category.
 Note that these methods do not employ additional architectural parameters for supernet training or search.
 They typically train the supernet by randomly sampling subnetworks and training them as depicted in Figure 1(a).
-The post-hoc blackbox search relies on using the performance of subnetworks sampled from the trained supernet as a proxy for true performance on the unseen test set.
-To contrast with this two-stage approach, we refer to traditional gradient-based NAS approaches as single stage methods.
-While to date, weight-entangled spaces have only been explored with two-stage methods, and cell-based spaces have only been optimized with single-stage approaches, in this paper we bridge the gap between these parallel sub-communities.
+The post-hoc black-box search relies on using the performance of subnetworks sampled from the trained supernet as a proxy for true performance on the unseen test set.
+To contrast with this two-stage approach, we refer to traditional gradient-based NAS approaches as single-stage methods.
+Figure 1: (a) Two-Stage NAS with WE (Algorithm 3): dotted paths show operation choices not sampled at the given step (b) Single-Stage NAS with WS (Algorithm 4): every operation choice is evaluated independently and contributes to the output feature map with corresponding architecture parameters (c) Single-Stage NAS with WE (Algorithm 1): operation choices superimposed with corresponding architecture parameters.
+The architecture parameters for the three operation choices are represented by ${[\alpha_{i}]}_{i=1}^{3}$ and ${[\beta_{i}]}_{i=1}^{3}$.
+The operation weights, or choices, are symbolized by cubes (for convolutional layers) or rectangles (for feedforward layers) in various colors.
+In scenarios (a) and (c), due to weight entanglement, the smaller weights are effectively structured subsets of the larger weights.
+Conversely, in (b), through weight-sharing, operation weights are maintained independently from each other.
+In both (b) and (c), to determine the optimal architecture, the operations associated with the highest architecture parameter value are selected.
+This selection process applies to the choice of kernel size and the output dimension of the feedforward network.
+To date, weight-entangled spaces have only been explored with two-stage methods, and cell-based spaces only with single-stage approaches.
+In this paper, we bridge the gap between these parallel sub-communities.
 We do so by addressing the challenges associated with integrating off-the-shelf single-stage NAS methods with weight-entangled search spaces.
 After a discussion of related work (Section 2), we make the following main contributions:
-  • We propose a generalized scheme to apply single-stage methods to weight-entangled spaces, while maintaining search efficiency and efficacy at larger scales (Section 3, with visualizations in Figure 1(c) and Figure 2).
+  1. 1.
+    We propose a generalized scheme to apply single-stage methods to weight-entangled spaces while maintaining search efficiency and efficacy at larger scales (Section 3, with visualizations in Figure 1(c) and Figure 2 and Figure 3).
 We refer to this method as TangleNAS.
-  • We propose a fair playground for the comparative evaluation of single and two-stage methods (Section 4.1) and study the effect of weight-entanglement in conventional cell-based search spaces (i.e., NASBench201 and the DARTS search space) (Section 4.2).
-  • We evaluate our proposed generalized scheme for single-stage methods across a diverse set of weight-entangled macro search spaces and tasks, from image classification (Section 4.3.1, Section 4.3.2) to language modeling (Section 4.3.3).
-  • We conduct a comprehensive evaluation of the properties of single and two-stage approaches including any-time-performance, memory consumption, robustness to training fraction and effect of fine-tuning (Section 5), demonstrating that our generalized gradient-based NAS method achieves the best of single and two-stage methods: the enhanced performance, improved supernet fine-tuning properties, superior any-time performance of single-stage methods, and the low memory consumption of two-stage methods .
+  2. 2.
+    We propose a unified evaluation framework for the comparative evaluation of single and two-stage methods (Section 4.1) and study the effect of weight-entanglement in conventional cell-based search spaces (i.e., NAS-Bench-201 and the DARTS search space) (Section 4.2).
+  3. 3.
+    We evaluate our proposed generalized scheme for single-stage methods across a diverse set of weight-entangled macro search spaces and tasks, from image classification (Section 4.3.1 and Section 4.3.2) to language modeling (Section 4.3.3).
+  4. 4.
+    We conduct a comprehensive evaluation of the properties of single and two-stage approaches (Section 5), demonstrating that our generalized gradient-based NAS method achieves the best of single and two-stage methods: the enhanced performance, improved supernet fine-tuning properties, superior any-time performance of single-stage methods, and low memory consumption of two-stage methods.
 To facilitate reproducibility, our code is openly accessible here.
 2 Related Work
-Weight-sharing was first introduced in ENAS (Pham et al., 2018), which reduced the computational cost of NAS by 1000$\times$ compared to previous methods.
-However, since this method used reinforcement learning its performance was quite brittle.
-Bender et al.
-(2018) simplified the technique, showing that searching for good architectures is possible by training the supernet directly with stochastic gradient descent.
-This was followed by DARTS (Liu et al., 2019), which set the cornerstone for efficient and effective gradient-based, single-stage, NAS approaches.
-DARTS, however, had prominent failure modes, such as its discretization gap and convergence towards parameter-free operations, as outlined in Robust-DARTS (Zela et al., 2020).
-Numerous gradient-based one-shot optimization techniques were developed since then (Cai et al., 2019; Nayman et al., 2019; Xu et al., 2019; Dong and Yang, 2019; Hu et al., 2020; Li et al., 2021; Chen et al., 2021b; Zhang et al., 2021).
-Among these, we highlight DrNAS (Chen et al., 2021b), which we will use in our experiments as a representative of gradient-based NAS methods.
+[…]
+DARTS, however, had prominent failure modes, such as its discretization gap and convergence towards parameter-free operations (white-arxiv23a), as outlined in Robust-DARTS (zela-iclr20a).
+Numerous gradient-based one-shot optimization techniques were developed since then (cai-iclr19a; nayman-neurips19a; wang-ijcai21; dong-cvpr19a; hu-cvpr20a; li-iclr21b; chen-iclr21a; zhang-icml21a).
+Amongst these, we highlight DrNAS (chen-iclr21a), which we will use in our experiments as a representative of gradient-based NAS methods.
 DrNAS treats one-shot search as a distribution learning problem, where the parameters of a Dirichlet distribution over architectural parameters are learned to identify promising regions of the search space.
-Despite the remarkable performance of single-stage methods, they are not directly applicable to some real-world architectural domains, such as transformers, because of the macro-level structure of these search spaces.
-DASH (Shen et al., 2022) employs a DARTS-like methodology to optimize CNN topologies (i.e.
-kernel size, dilation) for a diverse set of tasks, reducing computational complexity by appropriately padding and summing kernels with different sizes and dilations.
-FBNet-v2 (Wan et al., 2020) makes an attempt along these lines for CNN topologies, but its methodology is not easily extendable to search spaces like transformers with multiple interacting modalities, such as embedding dimension, number of heads, expansion ratio, and depth.
-Weight-entanglement, on the other hand, provides a more effective way of weight-sharing exclusive to macro-level architectural spaces.
-The concept of weight-entanglement was developed in slimmable networks (Yu et al., 2018; Yu and Huang, 2019), OFA (Cai et al., 2020) and BigNAS (Yu et al., 2020) in the context of convolutional networks (see also AtomNAS (Mei et al., 2019)) and later spelled out in AutoFormer (Chen et al., 2021a), where it was applied to the transformer architecture.
-Single-path-one-shot (SPOS) methods (Guo et al., 2020) have shown a lot of promise in searching weight-entangled spaces.
-SPOS trains a supernet by uniformly sampling single paths (one at a time to limit memory consumption) and then training the weights along that path.
+Despite the remarkable performance of single-stage methods, they are not directly applicable to some real-world architectural domains, such as transformers, due to the macro-level structure of these search spaces.
+DASH (shen-neurips22) employs a DARTS-like method to optimize CNN topologies (i.e., kernel size, dilation) for a diverse set of tasks, reducing computational complexity by appropriately padding and summing kernels with different sizes and dilations.
+FBNet-v2 (wan-cvpr20) and MergeNAS (wang-ijcai21) make an attempt along these lines for CNN topologies, but their methodology is not easily extendable to search spaces like transformers with multiple interacting modalities, such as embedding dimension, number of heads, expansion ratio, and depth.
+Weight-entanglement, on the other hand, provides a more effective way of weight-sharing, exclusive to macro-level architectural spaces.
+In weight-entanglement, operations with weights of smaller dimensionality are structured subsets of the largest dimension.
+Hence, the total number of parameters in the supernet is the same as the number of parameters in the largest architecture in the space.
+The concept of weight-entanglement was developed in slimmable networks (yu-iclr18; yu-cvpr19), OFA (cai-iclr20a) and BigNAS (yu-eccv20) in the context of convolutional networks (see also AtomNAS (mei-iclr19)) and later spelled out in AutoFormer (chen-iccv21b) and applied to the transformer architecture.
+Single-path-one-shot (SPOS) methods (guo-eccv20a) have shown a lot of promise in searching weight-entangled spaces.
+SPOS trains a supernet by uniformly sampling paths (one at a time to limit memory consumption) and then training the weights along that path.
 The supernet training is followed by a black-box search that uses the performance of the models sampled from the trained supernet as a proxy.
-OFA used a similar idea to optimize different dimensions of CNN architectures, such as its depth, width, kernel size and resolution.
-Additionally, it enforced training of larger to smaller sub-networks sequentially to prevent interference between sub-networks.
+OFA used a similar idea to optimize different dimensions of CNN architectures, such as its depth, width, kernel size, and resolution.
+Additionally, it enforced training of larger to smaller subnetworks sequentially to prevent interference between subnetworks.
 Subsequently, AutoFormer adopted the SPOS method to optimize a weight-entangled space of transformer architectures.
-In this work, we demonstrate how single-stage methods can be applied to macro-level search spaces with weight-entanglement, where one can benefit from the time efficiency and remarkable effectiveness of modern differentiable NAS optimizers, while concurrently maintaining memory efficiency of the weight-entangled space.
-While we chose to adopt DrNAS as the primary approach for our exploration of the weight-entangled space in this study, it is worth highlighting that our methodology is generally applicable to other gradient-based NAS methods.
+In this work, we demonstrate the application of single-stage methods to macro-level search spaces with weight-entanglement.
+This approach leverages the time efficiency and effectiveness of modern differentiable NAS optimizers, while maintaining the memory efficiency inherent to the weight-entangled space.
+Although DrNAS is our primary method for exploring weight-entangled spaces, our methodology is broadly applicable to other gradient-based NAS methods.
 3 Methodology: Single-Stage NAS with Weight-Superposition
+Figure 2: Weight superposition with architecture parameters ${\alpha_{i}}_{i=1}^{3}$ for kernel size search. Supernet weight matrix (LHS) is adapted to gradient-based methods (RHS).
+Computational Efficiency
 Our primary goal in this work is to effectively apply single-stage NAS to search spaces with macro-level architectural choices.
 To reduce the memory consumption and preserve computational efficiency, we propose two major modifications to single-stage methods.
-In our framework, the operators on any given edge inherit the weights of the largest operator on that edge.
+Firstly, the weights of the operations on every edge are shared with the corresponding weights of the largest operator on that edge.
 This reduces the size of the supernet to the size of the largest individual architecture in the search space.
-Figure 1: (a) Two-Stage NAS with WE (Algorithm 3): dotted paths show operation choices not sampled at the given step (b) Single-Stage NAS with WS (Algorithm 4): every operation choice is evaluated independently and contributes to the output feature map with corresponding architecture parameters (c) Single-Stage NAS with WE (Algorithm 1): operation choices superimposed with corresponding arhictecture parameters.
-Figure 2: Supernet weight matrix (LHS) adapted to gradient-based methods (RHS).
-All operation choices are zero-padded to match the dimension of the largest operation and superimposed with given architecture parameter.
-(a) Weight superposition for a single dimension choice (e.g.
-kernel size) with architecture parameter $\alpha_{i}$ (b) Weight superposition for combination of multiple dimensions (e.g.
-embedding dimension and expansion ratio) with architecture parameters $\alpha_{i}$, $\beta_{j}$.
-White areas inside the dashed gray boundaries indicate zero-padding and color values reflect scaling.
+Figure 3: Combi-superposition with parameters $\alpha_{i}\beta_{j}$. Supernet weight matrix (LHS) is adapted to gradient-based methods (RHS).
+Secondly, to compute operation mixture, we take each operation choice, zero-pad to match the size of the largest choice, and then sum these, with each one multiplied by its respective architectural parameter.
+The resulting operation thus represents a mixture of different choices on an edge.
+This approach, visualized in Figure 1(c), contrasts with single-stage NAS methods, which weigh the outputs of operations on a specific edge using architectural parameters (refer to Figure 1(b)).
+Figure 2 provides an overview of the idea for a single architectural choice, such as the kernel size.
+This is equivalent to taking the largest operation and re-scaling the weights of each sub-operation by corresponding architectural parameters followed by summation of the weights (see the right-most weight matrix in Figure 2).
 Algorithm 1 TangleNAS
@@
 $\sum_{i=1}^{N}f(\alpha_{i})=1$
-2:  $Cell_{j}\leftarrow DAG(\mathcal{O}_{j},{\mathcal{W}_{max}}_{j})$ /* defined for j=1...M */
-3:  $Supernet\leftarrow\cup_{j}^{M}Cell_{j}\cup\mathcal{A}$
+[…]
+[…]
 4:/* example of forward propagation on the cell */
@@
 11: $\mathcal{A}=\mathcal{A}-\gamma\nabla_{\mathcal{A}}\mathcal{L}_{val}({\mathcal{W}_{\max}}^{*},\mathcal{A})$
-12:   $\mathcal{\mathcal{missing}}{W}_{\max}=\mathcal{W}_{\max}-\eta\nabla_{\mathcal{W}_{\max}}\mathcal{L}_{train}(\mathcal{W}_{\max},\mathcal{A})$
+12: $\mathcal{\mathcal{}}{W}_{\max}=\mathcal{W}_{\max}-\eta\nabla_{\mathcal{W}_{\max}}\mathcal{L}_{train}(\mathcal{W}_{\max},\mathcal{A})$
 13: /* Architecture Selection */
 14: $selected\_arch\leftarrow$ $\arg\max(\mathcal{A})$
-Then, instead of weighting the outputs of operations on a given edge by architectural parameters, as single-stage NAS methods do (see Figure 1(b)), we simply take each operation choice, zero-pad that to match the dimensions of the largest operation and sum these terms, each weighted by the corresponding architectural parameter.
-Figure 2(a) provides an overview of the idea for a single architectural choice such as the kernel size.
-This is equivalent to taking the largest operation and re-scaling the weights of each sub-operation by the sum of architectural parameters of operations it is embedded in (see right-most weight matrix in Figure 2(a).
-This process results in a superposition of operation weights whose structural property is similar to the largest operation.
-Consequently, a single forward pass suffices to capture the effect of all operational choices, thus making our approach computationally efficient.
-Furthermore, to accommodate for operations depending on two or more architectural dimension choices we introduce the combi-superposition outlined in Figure 2(b) and Algorithm 2.
-This allows us to apply any arbitrary gradient-based NAS method, such as DARTS, GDAS or DrNAS, to macro-level search spaces that leverage weight-entanglement, without additional memory consumption and computational cost during forward propagation.
-We refer to our memory and compute optimal single-stage architecture search method as TangleNAS and provide an overview of the approach in Algorithm 1.
+Weight Superposition
+Weight Superposition is defined as a weighted summation of subsets of the largest weight matrix, to obtain weights with structural properties comparable to the largest operation.
+Consequently, a single forward pass on the superimposed weights suffices to capture the effect of all operational choices, thus making our approach computationally efficient.
+See Figure 2 for details on weight superposition of the kernel size.
+Combi-Superposition
+An operation may depend on two or more architectural decisions.
+Consider, for example, embedding dimension and intermediate MLP dimension (where the latter depends on the former by a searchable multiplicative factor i.e., the MLP ratio).
+To accommodate this, we introduce combi-superposition outlined in Figure 3 and in Algorithm 2.
+Combi-superposition superimposes the weights across multiple different architectural dimensions, allowing for search across different interacting architectural modalities.
+Algorithm
+These concepts allow us to apply any arbitrary gradient-based NAS method, such as DARTS, GDAS, or DrNAS, to macro-level search spaces that leverage weight-entanglement without incurring additional memory and computational costs during forward propagation.
+We name this single-stage architecture search method TangleNAS.
+See Algorithm 1 for an overview of the bi-level optimization framework with weight-superposition.
 The operation $f$ in Algorithm 1 determines the differentiable optimizer used in the method.
-A $softmax$ function corresponds to DARTS, while sampling from the Dirichlet distribution corresponds to DrNAS.
-In all experiments of the upcoming sections we use DrNAS as a basic gradient-based NAS method.
-We therefore simply refer to this combination as TangleNAS throughout the paper.
+For example, $f$ is a softmax function for DARTS and a function that samples from the Dirichlet distribution for DrNAS.
+We use DrNAS as the primary gradient-based NAS method in all our experiments, and refer to DrNAS used in conjunction with weight-entangled spaces as TangleNAS in the remainder of the paper.
 4 Experiments
@@
 4.1 Toy search spaces
-We begin the evaluation of TangleNAS on two relatively compact toy search spaces that we designed as a contribution to the community to allow faster iterations of algorithm development:
-  • Toy cell space: a small version of the DARTS space; architectures are evaluated on the Fashion-MNIST dataset.
-  • Toy conv-macro space: a small space inspired by MobileNet, including kernel sizes and the number of channels in each convolution layer; architectures are evaluated on CIFAR-10.
-We describe these spaces in Appendix C, including links to code for these open source toy benchmarks.
+We begin the evaluation of TangleNAS on two compact toy search spaces that we designed as a contribution to the community to allow faster iterations of algorithm development:
+  • Toy cell space: a small version of the DARTS space; architectures are evaluated on the Fashion-MNIST dataset (xiao2017fashion).
+  • Toy conv-macro space: a small space inspired by MobileNet, including kernel sizes and the number of channels in each convolution layer as architectural decisions.
+The architectures are evaluated on CIFAR-10.
+We describe these spaces in Appendix D, including links to code for these open source toy benchmarks.
 The results of these experiments are summarized in Tables 2 and 2.
-Across both of these search spaces, TangleNAS outperforms its two-stage counterparts.
-Also, DrNAS without weight entanglement performs very poorly on the macro level search space.
-Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset
-Search Type Optimizer Supernet type Test acc Search Time(hrs)
-Single-Stage DrNAS WS $91.190\pm 0.049$ 6.277
-TangleNAS WE $\mathbf{91.300\pm 0.023}$ 6.222
-Two-Stage SPOS+RS WE $90.680\pm 0.253$ 15.611
-SPOS+ES WE $90.317\pm 0.223$ 13.244
+In both of these search spaces, TangleNAS outperforms its two-stage counterparts over 4 seeds.
+Additionally, DrNAS without weight-entanglement performs extremely poorly on the macro level search space (equivalent to a random classifier).
+Search Type Optimizer Supernet type Test acc (%) Search Time (hrs)
+Single-Stage DrNAS WS $91.190\pm 0.049$ 6.3
+TangleNAS WE $\mathbf{91.300\pm 0.023}$ 6.2
+Two-Stage SPOS+RS WE $90.680\pm 0.253$ 15.6
+WE $90.317\pm 0.223$ 13.2
 Optimum - - 91.630 -
-Search Type Optimizer Supernet type Test acc Search Time(hrs)
-Single-Stage DrNAS WS $10\pm 0.00$ 12.361
-TangleNAS WE $\mathbf{82.495\pm.0.461}$ 8.55
-Two-Stage SPOS+RS WE $81.253\pm 0.6717$ 21.676
-SPOS+ES WE $81.890\pm 0.800$ 26.359
+Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset.
+Search Type Optimizer Supernet type Test acc (%) Search Time (hrs)
+Single-Stage DrNAS WS $10\pm 0.000$ 12.4
+TangleNAS WE $\mathbf{82.495\pm 0.461}$ 8.6
+Two-Stage SPOS+RS WE $81.253\pm 0.672$ 21.7
+WE $81.890\pm 0.800$ 26.4
 Optimum - - 84.410 -
-Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset
-Table 2: Evaluation on the toy conv-macro search space on the CIFAR10 dataset. Note that DrNAS with weight sharing converges to a degenerate architecture here.
+Table 2: Evaluation on the toy conv-macro search space on the CIFAR-10 dataset.
 4.2 Cell-based search spaces
-We now start our comparative analysis of single and two-stage approaches by applying them to cell-based spaces that serve as focal points in the single-stage NAS literature to evaluate TangleNAS against DrNAS and SPOS.
-Here, we use the popular NAS-Bench-201 (NB201) (Dong and Yang, 2020) and DARTS (Liu et al., 2019) search spaces.
-We refer the reader to Appendix C for details about these spaces and Section E.4 for the experimental setup.
+We now begin our comparative analysis of single and two-stage approaches by applying them to cell-based spaces, which are central in the single-stage NAS literature.
+We evaluate TangleNAS against DrNAS and SPOS on these spaces.
+Here, we use the widely studied NAS-Bench-201 (NB201) (dong2020bench) and DARTS (liu-iclr19a) search spaces.
+We refer the reader to Appendix D for details about these spaces and Appendix F.4 for the experimental setup.
+We evaluate each method with 4 different random seeds.
+Search Type Optimizer Supernet CIFAR-10 (%) ImageNet (%) Search Time (hrs)
+Single-Stage DrNAS WS $2.625\pm 0.075$ 26.290 9.1
+TangleNAS WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.691}$ 7.4
+Two-Stage SPOS+RS WE $2.965\pm 0.072$ 27.114 18.7
+SPOS+ES WE $3.200\pm 0.065$ 27.320 14.8
+Table 3: Comparison of test errors of single and two-stage methods on the DARTS search space.
 Our contribution on these spaces is two-fold.
-First, we study the effects of weight-entanglement on cell-based spaces in conjunction with single-stage methods.
+Firstly, we study the effects of weight-entanglement on cell-based spaces in conjunction with single-stage methods.
 To this end, we entangle the weights of similar operations with different kernel sizes on both search spaces.
-For NB201, the weights of the 1x1 and 3x3 convolutions are entangled, and in the DARTS search space, the weights of dilated and separable convolutions with kernel sizes 3x3 and 5x5 are entangled.
-Second, we study SPOS on the NB201 and DARTS search spaces.
+For NB201, the weights of the $\times1$ and $\times3$ convolutions are entangled, and in the DARTS search space, the weights of dilated and separable convolutions with kernel sizes $\times3$ and $\times5$ are (separately) entangled.
+Secondly, we study SPOS on the NB201 and DARTS search spaces.
 To the best of our knowledge, we are the first to study a two-stage method like SPOS in such cell search spaces.
@@
 We note that overall, the SPOS methods are ineffective in these cell search spaces.
-Search Type Optimizer Supernet CIFAR10 CIFAR100 ImageNet16-120 Search Time (hrs)
-Single-Stage DrNAS WS $\mathbf{94.36\pm 0.00}$ $72.245\pm 0.732$ $\mathbf{46.37\pm 0.00}$ 20.9166
-TangleNAS WE $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$ 20.3611
-Two-Stage SPOS+RS WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.1460$ 29.51388
-SPOS+ES WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.0149$ 26.74721
-Table 3: Comparison of test-accuracies of single and two-stage methods on NB201 search space
-Search Type Optimizer Supernet CIFAR10 ImageNet Search Time(hrs)
-Single-Stage DrNAS WS $2.625\pm 0.075$ 26.29 9.12
-TangleNAS WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.691}$ 7.44
-Two-Stage SPOS+RS WE $2.965\pm 0.072$ 27.114 18.7444
-SPOS+ES WE $3.200\pm 0.065$ 27.320 14.794
-Table 4: Comparison of test errors of single and two-stage methods on the DARTS search space
+Search Type Optimizer Supernet CIFAR-10 CIFAR-100 ImageNet16-120 Search Time (hrs)
+Single-Stage DrNAS WS $\mathbf{94.360\pm 0.000}$ $72.245\pm 0.732$ $\mathbf{46.370\pm 0.000}$ 20.9
+TangleNAS WE $\mathbf{94.360\pm 0.000}$ $\mathbf{73.510\pm 0.000}$ $\mathbf{46.370\pm 0.000}$ 20.4
+Two-Stage SPOS+RS WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.146$ 29.5
+SPOS+ES WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.015$ 26.7
+Table 4: Comparison of test-accuracies of single and two-stage methods on NB201 search space.
 4.3 Macro Search Spaces
 Given the promising results of TangleNAS on the toy and cell-based spaces, we now extend our evaluation to the home base of two-stage methods.
-We study TangleNAS on a vision transformer space (AutoFormer-T) and a convolutional space (MobileNetV3), which have been proposed and explored using two-stage methods by Chen et al.
-(2021a) and Cai et al.
-(2020), respectively, as well as a language model transformer search space built around GPT-2 Radford et al.
-(2019).
+We study TangleNAS on a vision transformer space (AutoFormer-T and -S) and a convolutional space (MobileNetV3), both of which were previously proposed and examined using two-stage methods by chen-iccv21b and cai-iclr20a, respectively.
+Additionally, we explore a language model transformer search space centered around GPT-2 (radford-openaiblog19a).
 4.3.1 AutoFormer
-Search Type Optimizer CIFAR10 CIFAR100
-Inherit Fine-tune Retrain Params FLOPS Inherit Fine-tune Retrain Params FLOPS
-Single-Stage TangleNAS 93.57 $\mathbf{97.702\pm 0.017}$ $\mathbf{97.872\pm 0.054}$ 8.6847M 1.946 76.59 $\mathbf{82.615\pm 0.064}$ $\mathbf{82.668\pm 0.161}$ 8.64874M 1.9388G
-Two-Stage SPOS+RS 94.29 $97.605\pm 0.038$ $97.767\pm 0.024$ 8.51161M 1.91G 78.210 $82.407\pm 0.026$ $82.210\pm 0.14242$ 8.47558M 1.9046G
-SPOS+ES 94.10 $97.632\pm 0.047$ $97.6425\pm 0.023$ 7.230286M 1.6595 77.97 $82.517\pm 0.140$ $82.5175\pm 0.114$ 8.2447M 1.859G
-Table 5: Evaluation on the AutoFormer-T space for CIFAR10 and CIFAR100.
-We evaluate TangleNAS on the AutoFormer-T space introduced by Chen et al.
-(2021a), which is based on vision transformers.
+Search Type Optimizer CIFAR-10 CIFAR-100
+Inherit (%) Fine-tune (%) Retrain (%) Params FLOPS Inherit (%) Fine-tune (%) Retrain (%) Params FLOPS
+Single-Stage TangleNAS 93.570 $\mathbf{97.702\pm 0.017}$ $\mathbf{97.872\pm 0.054}$ 8.685M 1.946G 76.590 $\mathbf{82.615\pm 0.064}$ $\mathbf{82.668\pm 0.161}$ 8.649M 1.939G
+Two-Stage SPOS+RS 94.290 $97.605\pm 0.038$ $97.767\pm 0.024$ 8.512M 1.910G 78.210 $82.407\pm 0.026$ $82.210\pm 0.142$ 8.476M 1.905G
+SPOS+ES 94.100 $97.632\pm 0.047$ $97.643\pm 0.023$ 7.230M 1.659G 77.970 $82.517\pm 0.140$ $82.518\pm 0.114$ 8.245M 1.859G
+Table 5: Test Accuracies on the AutoFormer-T space for CIFAR-10 and CIFAR-100 (across 4 random seeds)
+We evaluate TangleNAS on the AutoFormer-T and AutoFormer-S spaces introduced by chen-iccv21b, based on vision transformers.
 The search space consists of the choices of embedding dimensions and number of layers, and for each layer, its MLP expansion ratio and the number of heads it uses to compute attention.
 More details can be found in Table 20 in the appendix.
-The embedding dimension choice is held constant throughout the network while the number of heads and the MLP expansion ratio varies for every layer.
-This amounts to a total search space size of about $10^{13}$ architectures.
-We train our supernet using the same training hyperparameters as used in AutoFormer.
-NAS Method SuperNet-Type ImageNet Datasets Params FLOPS
-CIFAR10 CIFAR100 Flowers Pets Cars
-SPOS+ES AutoFormer-T 75.474 98.019 86.369 98.066 91.558 91.935 5.893M 1.396G
-TangleNAS AutoFormer-T 78.842 98.249 88.290 98.066 92.347 92.396 8.98108M 2.00G
-SPOS+ES AutoFormer-S 81.700 99.10 90.459 97.90 94.8529 92.5447 22.9M 5.1G
-TangleNAS AutoFormer-S 81.964 99.12 90.459 98.3257 95.07 92.3707 28.806M 6.019G
+The embedding dimension remains constant across the network, while the number of heads and the MLP expansion ratio change for each layer.
+This results in a search space of about $10^{13}$ architectures.
+We train our supernet using the same training hyperparameters and pipeline as used in AutoFormer, and use its evolutionary search as our baseline.
+SuperNet-Type NAS Method ImageNet (%) Datasets Params FLOPS
+CIFAR-10 (%) CIFAR-100 (%) Flowers (%) Pets(%) Cars(%)
+AutoFormer-T SPOS+ES 75.474 98.019 86.369 98.066 91.558 91.935 5.893M 1.396G
+AutoFormer-T TangleNAS 78.842 98.249 88.290 98.066 92.347 92.396 8.981M 2.000G
+AutoFormer-S SPOS+ES 81.700 99.100 90.459 97.900 94.853 92.545 22.900M 5.100G
+AutoFormer-S TangleNAS 81.964 99.120 90.459 98.326 95.070 92.371 28.806M 6.019G
 Table 6: Evaluation on the AutoFormer-T space on downstream tasks.
-ImageNet accuracies are obtained through inheritance, whereas accuracies for the other datasets are achieved through fine-tuning the imagenet-pretrained model.
-Table 5 and Table 6 evaluates TangleNAS against AutoFormer on the AutoFormer-T space.
+ImageNet-1k validation accuracies are obtained through inheritance, whereas the test accuracies for the other datasets are achieved through fine-tuning the ImageNet-pretrained model.
+In Tables 5 and 6, we evaluate TangleNAS against AutoFormer on the AutoFormer-T and AutoFormer-S spaces.
 Interestingly, we observe that although AutoFormer sometimes outperforms TangleNAS upon inheritance from the supernet, the TangleNAS architectures are always better upon fine-tuning and much better upon retraining.
-For ImageNet-1k we obtain an accuracy of $78.842$% by inheriting the weights of the optimal architecture in contrast to $75.474$% by unconstrained evolutionary search on the on AutoFormer-T space (Chen et al., 2021a), i.e.
-a net improvement of $3.368\%$ (see Table 6).
+For ImageNet-1k we obtain an improvement of $3.368\%$ and $0.264\%$ on AutoFormer-T and AutoFormer-S spaces, respectively (see Table 6).
 4.3.2 MobileNetV3
-Search Type Optimizer Top-1 acc Params FLOPS
-Single-Stage TangleNAS 77.424 7.58M 528.80M
-Two-Stage OFA+RS 77.046 6.87M 369.16M
-OFA+ES 77.050 7.21M 420.50M
-Largest Arch - 77.336 7.66M 566.17M
-Table 7: Evaluation on MobileNetV3
 Next, we study a convolutional search space based on the MobileNetV3 architecture.
 The search space is defined in Table 19 in the appendix and contains about $2\times 10^{19}$ architectures.
-This follows from the search space designed by OFA (Cai et al., 2020), which searches for kernel-size, number of blocks and channel-expansion factor.
-Unlike OFA, we do not use the progressive shrinking scheme for the supernet training but activate all choices in our supernet at all times.
-Table 7 shows that on this OFA search space, TangleNAS still outperforms OFA (based on both unconstrained evolutionary and random search on the pre-trained OFA supernet).
+This follows from the search space designed by OFA (cai-iclr20a), which searches for kernel size, number of blocks, and channel-expansion factor.
+Search Type Optimizer Top-1 acc (%) Params FLOPS
+Single-Stage TangleNAS 77.424 7.580M 528.80M
+Two-Stage OFA+RS 77.046 6.870M 369.160M
+OFA+ES 77.050 7.210M 420.500M
+Largest Arch - 77.336 7.660M 566.170M
+Table 7: Evaluation on MobileNetV3.
+During the supernet training, we activate all choices in our supernet at all times.
+Table 7 shows that on this OFA search space, TangleNAS outperforms OFA (based on both unconstrained evolutionary and random search on the pre-trained OFA supernet).
 Notably, TangleNAS even yields an architecture with higher accuracy than the largest architecture in the space (while OFA yields worse architectures).
-4.3.3 Language Modelling Space
-Finally, given the growing interest in efficiently training large language models and recent developments in scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), we study our efficient and scalable TangleNAS method on a language-model space.
-We create our language model space around a smaller version of Andrej Karpathy’s nanoGPT model [^1] https://github.com/karpathy/nanoGPT.
-In this transformer search space, we search for the embedding dimension, number of heads, number of layers and MLP ratios as defined in Table 19 (in the appendix).
-For each of these transformer search dimensions we consider 3 choices.
-We evaluate our small language model on the TinyStories (Eldan and Li, 2023) dataset.
-We improve over the SPOS+ES and SPOS+RS baselines as presented in Table 8.
+4.3.3 Language Modeling (LM) Space
+Finally, given the growing interest in efficient large language models and recent developments in scaling laws (kaplan2020scaling; hoffmann2022training), we study our efficient and scalable TangleNAS method on a language-model space for two different scales.
+We create our language model space around a smaller version of nanoGPT model [^1] https://github.com/karpathy/nanoGPT and the model at its original size.
+In this transformer search space, we search (with 4 random seeds) for the embedding dimension, number of heads, number of layers, and MLP ratios (as defined in Table 18 of the appendix).
+We weight-entangle all of these by combi-superposition in four dimensions.
+Architecture Search-Type Loss Perplexity $\mathbf{\downarrow}$ Params Inference time (s)
+GPT-2 Manual 3.077 21.690 123.590M 113.30
+TangleNAS Automated 2.904 18.243 116.519M 102.50
+(a) Comparison of fine-tuning architecture discovered by TangleNAS and GPT-2 on Shakespeare dataset.
+Search Type Optimizer Test loss Perplexity $\mathbf{\downarrow}$ Params Inference Time (s)
+Single-Stage TangleNAS $\mathbf{1.412\pm 0.011}$ 4.104 87.010M 93.88
+Two-Stage SPOS+RS $1.433\pm 0.005$ 4.191 77.570M 85.17
+SPOS+ES $1.444\pm 0.013$ 4.238 78.750M 87.10
+(b) Comparison of single and two-stage methods on language model search space. We report the test loss and perplexity on the TinyStories dataset.
+Table 8: Evaluation on Language Model Space
+For each of these transformer search dimensions, we consider 3 different choices.We train our language models on the TinyStories (eldan-arxiv2023) dataset (for the smaller version of nanoGPT) and on OpenWebText (Gokaslan2019OpenWeb) (for nanoGPT at its original size).
+Furthermore, we fine-tune the model trained on OpenWebText on the Shakespeare dataset.
+On the smaller scale, we beat the SPOS+ES and SPOS+RS baselines as presented in Table 8(b).
 This improvement is statistically very significant with a two-tailed p-value of $0.0064$ for ours v/s SPOS+ES and p-value of $0.0127$ for ours v/s SPOS+RS.
-[…]
-Single-Stage TangleNAS $\mathbf{1.412\pm 0.011}$ 4.104 87.01M 93.88s
-Two-Stage SPOS+RS $1.433\pm 0.005$ 4.191 77.57M 85.169s
-SPOS+ES $1.444\pm 0.013$ 4.238 78.75M 87.10s
-Table 8: Comparison of single and two stage methods on language model search space. We report the test loss and perplexity on the tinystories dataset
+As shown in Table 8(a), we obtain a smaller model on OpenWebText while achieving better perplexity after fine-tuning on Shakespeare [^2] https://huggingface.co/datasets/karpathy/tiny_shakespeare.
+This model is also very efficient during inference time in comparison to GPT-2.
 5 Results and Discussion
-For a more in-depth evaluation, we now compare different properties of single-stage and two-stage methods, starting from their any-time performance, impact of the portion of train-val split on their performance and centered-kernel alignment of the supernet feature maps.
-Moreover, we also study the effect of pretraining, fine-tuning and retraining on the AutoFormer space for CIFAR10 and CIFAR100 and downstream performance of imagenet pre-trained best model on several classification datasets.
-Finally, we discuss the insights derived from our single-stage approach in designing architectures on real world tasks.
+For a more thorough evaluation, we now compare different properties of single and two-stage methods, focusing on their any-time performance, the impact of the train-validation split ratios, and the Centered Kernel Alignment (kornblith2019similarity) (see Appendix H) of the supernet feature maps.
+Additionally, we study the effect of pretraining, fine-tuning and retraining on the AutoFormer space for CIFAR-10 and CIFAR-100, as well as the downstream performance of the best model pre-trained on ImageNet across various classification datasets.
+We conclude by discussing the insights derived from TangleNAS in designing architectures on real world tasks.
 Space and time complexity
-In practice, we observe that vanilla gradient-based NAS methods are memory and compute expensive in comparison to both two-stage methods and our TangleNAS approach with weight-superposition (described in Figure 2).
-In particular on the NB201 and DARTS search spaces we observe a 25.276% and 35.545% reduction in memory requirements for TangleNAS over DrNAS with WS.
-This issue only exacerbates for weight-entangled spaces like AutoFormer and MobileNetV3, making the application of vanilla gradient-based methods practically infeasible.
-This is because vanilla gradient-based methods incur an $\mathcal{O}(n)$ overhead over TangleNAS in terms of space and time complexity, where n is the number of operation choices.
-Any-Time performance
-An essential characteristic of any neural architecture search (NAS) method is the attainment of strong any-time performance.
-End users of NAS techniques frequently prioritize the expedited identification of competitive architectures within limited time constraints.
-This becomes particularly crucial in light of the escalating computational expenses associated with training sophisticated and large neural networks, such as Transformers.
-Strong any-time performance is a pivotal criterion for the effective applicability of NAS in such resource-intensive domains.
-We thus compare the any-time performance of TangleNAS.
-Figure 3 demonstrates that TangleNAS (DrNAS+WE) is faster than its baseline method DrNAS+WS.
+In practice, we observe that vanilla gradient-based NAS methods are memory and compute expensive in comparison to both two-stage methods and our TangleNAS approach with weight-superposition.
+While the time and space complexity of single-stage methods is $\mathcal{O}(n)$, where $n$ is the number of operation choices, TangleNAS, similar to two-stage methods, maintains $\mathcal{O}(1)$.
+On the NB201 and DARTS search spaces we observe a 25.28% and 35.54% reduction in memory requirements for TangleNAS over DrNAS with WS.
+This issue only exacerbates for weight-entangled spaces like AutoFormer, MobileNetV3 and GPT, making the application of vanilla gradient-based methods practically infeasible.
+Anytime performance
+NAS practitioners often emphasize rapid discovery of competitive architectures.
+This is especially important given the rising costs of training large and complex neural networks, like Transformers.
+Thus, strong anytime performance is crucial for the practical deployment of NAS in resource-intensive environments.
+Therefore, we examine the anytime performance of TangleNAS.
+Figure 4 demonstrates that TangleNAS (DrNAS+WE) is faster than its baseline method (DrNAS+WS).
 Similarly, Figure 5 shows that TangleNAS has better anytime performance than AutoFormer.
-(a) CIFAR100
+(a) CIFAR-100
 (b) ImageNet16-120
-Figure 3: Test accuracy evolution over epochs for NB201
-(a) CIFAR10
-(b) CIFAR100
-Figure 4: Any Time performance curves of AutoFormer vs. Ours
-Figure 5: MLP ratio trajectory for LM. Number of layers range from 1-12 and mlp ratio choice can be 2,3 or 4.
+Figure 4: Test accuracy evolution over epochs for NB201.
+(a) CIFAR-10
+(b) CIFAR-100
+Figure 5: Any Time performance curves of AutoFormer vs. Ours.
 Effect of fraction of training data
 One-shot NAS commonly employs a 50%-50% train-valid split for cell-based spaces and 80%-20% for weight-entangled spaces.
-To avoid potential confounding factors, we also study our method across different data splits for each search space.
-The results reported in Appendix A are largely similar across splits.
-In particular, we observe that single-stage methods are quite robust and performant across training fractions and search spaces, in comparison to two-stage methods.
+To eliminate possible biases, we tested our method on various data splits within each search space.
+The findings, detailed in Section B, show consistent results across these splits.
+Specifically, single-stage methods prove to be robust and performant across different training fractions and search spaces, compared to two-stage methods.
+Figure 6: MLP ratio trajectory for LM. Number of layers range from 1-12 and MLP ratio choice can be 2, 3 or 4.
+5.1 Insights from NAS
 Architecture design insights
-In transformer spaces reducing the mlp-ratio in the initial layers has a relatively low impact on performance (Figure 5) and can often work competitively or outperform handcrafted architectures.
+In transformer spaces, reducing the MLP-ratio in the initial layers has a relatively low impact on performance (Figure 6) and can often work competitively or outperform handcrafted architectures.
 This observation is consistent across ViT and Language Model spaces.
-Conversely, number of heads and embedding dimension have a significant impact on the quantitative metric of choice.
+Conversely, number of heads and embedding dimension have a significant impact.
 Pruning a few of the final layers also has a relatively low impact on performance.
-In the MobileNetV3 space, we find a strong preference for larger number of channels and larger network depth.
-In contrast, we discover that scaling laws for transformers may not necessary apply in convolutional spaces especially for kernel sizes.
-We observe that earlier layers favor 5x5 kernel sizes while later ones prefer 7x7 kernel sizes (3 being the smallest and 7 the largest).
+In the MobileNetV3 space, we find a strong preference for a larger number of channels and larger network depth.
+In contrast, we discover that scaling laws for transformers may not necessarily apply in convolutional spaces, especially for kernel sizes - the earlier layers favor $\times5$ kernel sizes while later ones prefer $\times7$ (3 being the smallest and 7 the largest).
 Effect of pretraining, fine-tuning and retraining
-We study the impact of inheriting, fine-tuning, and retraining on the CIFAR10 and CIFAR100 datasets on the AutoFormer space.
-We observe that retraining almost always outperforms fine-tuning and inheriting.
-This brings to question the rank correlation between the inherited accuracy and retraining accuracy of two-stage methods and the training interference (Xu et al., 2022) in two-stage methods.
-Strong rank-correlation is desirable especially in two-stage methods as these methods rely on the performance proxy upon inheritance for the black-box search.
-Furthermore, we observe that while the SPOS+RS and SPOS+ES methods are performant upon inheritance, TangleNAS outperforms these models after fine-tuning and training from scratch.
-Improved supernet training properties for single-stage methods is further supported by Centered Kernel Alignment (CKA) analysis (Kornblith et al., 2019) (Table 25).
+We examined the effects of inheriting, fine-tuning, and retraining in the AutoFormer space on the CIFAR-10 and CIFAR-100 datasets.
+We observe that retraining generally surpasses both fine-tuning and inheriting.
+This raises questions about the correlation between inherited and retrained accuracies of architectures in two-stage methods and the potential training interference identified by xu2022analyzing.
+A strong correlation is crucial for two-stage methods, which use the performance of architectures with inherited weights as a proxy for true performance in the black-box search.
+Indeed, while the SPOS+RS and SPOS+ES methods perform well with inherited weights, TangleNAS exceeds their performance after fine-tuning and retraining.
 ImageNet-1k pre-trained architecture on downstream tasks
-Finally, we study the impact on fine-tuning the best model obtained from the search on downstream datasets.
-We follow the fine-tuning pipeline proposed in AutoFormer and fine-tune on different fine and coarse-grained datasets.
+Lastly, we study the impact on fine-tuning the best model obtained from the search on downstream datasets.
+We follow the fine-tuning pipeline proposed in AutoFormer and fine-tune on different fine- and coarse-grained datasets.
 We observe from Table 6 that the architecture discovered by TangleNAS on ImageNet is much more performant in fine-tuning to various datasets (CIFAR-10, CIFAR-100, Flowers, Pets and Cars) than the architecture discovered by SPOS.
-6 Conclusion
-In this paper, we investigate two of the most widely used NAS methods: single-stage and two-stage, each traditionally applied to different types of search spaces.
-We extend the applicability of single-stage NAS methods to weight-entangled spaces, a domain primarily explored by two-stage methods.
-By doing so, we aim to bridge the gap between these two distinct sub-fields.
-We empirically evaluate our proposed single-stage method, TangleNAS, on a diverse set of weight-entangled search spaces and tasks, showcasing its ability to outperform conventional two-stage NAS methods while enhancing search efficiency.
-Furthermore, we highlight the advantages of introducing weight-entanglement into traditional cell-based spaces, whenever feasible.
-Doing so not only improves the speed of single-stage methods but also reduces memory consumption.
-The modular nature of TangleNAS enables it to accommodate any differentiable optimizer and hence take advantage of the continually-improving set of methods for gradient-based NAS.
-Our results on various macro-level search spaces are already very encouraging and we therefore hope that this line of work will enable the NAS community to further improve modern architectures like Transformers.
-6.0.1 Acknowledgments
-This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the German Federal Ministry of Education and Research (BMBF, grant RenormalizedFlows 01IS19077C); the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no.
+6 Conclusion and Broader Impact
+In this paper, we compare single-stage and two-stage NAS methods, traditionally used for different search spaces, and introduce single-stage NAS to weight-entangled spaces, usually the domain of two-stage methods.
+We empirically evaluate our single-stage method, TangleNAS, on a diverse set of weight-entangled search spaces and tasks, showcasing its ability to outperform conventional two-stage NAS methods while enhancing search efficiency.
+Our positive results on macro-level search spaces suggest this approach could advance the development of modern architectures like Transformers within the NAS community.
+A recent work (klein2023structural) starts training of the supernet from the largest pretrained model, subsequently fine-tuning it.
+Our method, which now renders single-stage methods applicable to broader families of search spaces (e.g., transformers), can similarly benefit from initialization with pretrained models, achieving additional computational savings.
+We leave this for future work.
+This study addresses the high energy consumption associated with neural architecture search, particularly in black-box techniques that demand extensive computational resources to train many architectures from scratch.
+Our proposed method falls in the family of one-shot NAS methods, and hence significantly reduces energy usage and identifies efficient architectures with far less energy consumption than manual tuning.
+Acknowledgments
+This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no.
 101045765).
 Robert Bosch GmbH is acknowledged for financial support.
+The authors acknowledge support from ELLIS and ELIZA.
 Funded by the European Union.
@@
 Neither the European Union nor the ERC can be held responsible for them.
-Appendix A Training across data fractions
-Search Type Optimizer Train portion CIFAR10 CIFAR100
-Single-Stage TangleNAS 50% $97.715\pm 0.0877$ $82.5375\pm 0.11764$
-80% $\mathbf{97.872\pm 0.05369}$ $\mathbf{82.6675\pm 0.16073}$
-Two-Stage SPOS+RS 50% $97.6799\pm 0.0262$ $82.5372\pm 0.27954$
-80% $97.7674\pm 0.0239278$ $82.210\pm 0.14242$
-SPOS+RE 50% $97.77\pm 0.03752$ $82.354\pm 0.11974$
-80% $97.6425\pm 0.02286$ $82.5175\pm 0.1135$
-Table 9: Evaluation on the AutoFormer-T space for CIFAR10 and CIFAR100
+Appendix A Submission Checklist
+  1. 1.
+    For all authors…
+    1. (a)
+      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
+    2. (b)
+      Did you describe the limitations of your work? Yes, in Appendix I
+    3. (c)
+      Did you discuss any potential negative societal impacts of your work? NA
+    4. (d)
+      Did you read the ethics review guidelines and ensure that your paper conforms to them? https://2022.automl.cc/ethics-accessibility/ Yes
+  2. 2.
+    If you ran experiments…
+    1. (a)
+      Did you use the same evaluation protocol for all methods being compared (e.g., same benchmarks, data (sub)sets, available resources)? Yes
+    2. (b)
+      Did you specify all the necessary details of your evaluation (e.g., data splits, pre-processing, search spaces, hyperparameter tuning)? Yes
+    3. (c)
+      Did you repeat your experiments (e.g., across multiple random seeds or splits) to account for the impact of randomness in your methods or data? Yes, for experiments on smaller scales
+    4. (d)
+      Did you report the uncertainty of your results (e.g., the variance across random seeds or splits)? Yes, for experiments on smaller scales
+    5. (e)
+      Did you report the statistical significance of your results? Yes, we perform a two tailed p-value test on the tinystories dataset for language modeling
+    6. (f)
+      Did you use tabular or surrogate benchmarks for in-depth evaluations? Yes, We used NAS-Bench-201 which is tabular
+    7. (g)
+      Did you compare performance over time and describe how you selected the maximum duration? Yes, we study performance over compute budgets
+    8. (h)
+      Did you include the total amount of compute and the type of resources used (e.g., type of gpus, internal cluster, or cloud provider)? Yes
+    9. (i)
+      Did you run ablation studies to assess the impact of different components of your approach? No, We only add a single main component in our approach
+  3. 3.
+    With respect to the code used to obtain your results…
+    1. (a)
+      Did you include the code, data, and instructions needed to reproduce the main experimental results, including all requirements (e.g., requirements.txt with explicit versions), random seeds, an instructive README with installation, and execution commands (either in the supplemental material or as a url)?
+Yes
+    2. (b)
+      Did you include a minimal example to replicate results on a small subset of the experiments or on toy data? Yes
+    3. (c)
+      Did you ensure sufficient code quality and documentation so that someone else can execute and understand your code? Yes
+    4. (d)
+      Did you include the raw results of running your experiments with the given code, data, and instructions?
+Yes, we include the smaller pickle files for toy benchmarks we create and results logs wherever possible.
+We will share larger raw files eg: checkpoints upon acceptance, as anonymization of these resources is difficult.
+    5. (e)
+      Did you include the code, additional data, and instructions needed to generate the figures and tables in your paper based on the raw results? Yes
+  4. 4.
+    If you used existing assets (e.g., code, data, models)…
+    1. (a)
+      Did you cite the creators of used assets? Yes
+    2. (b)
+      Did you discuss whether and how consent was obtained from people whose data you’re using/curating if the license requires it? We do not use personal data
+    3. (c)
+      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? We do not use personal data
+  5. 5.
+    If you created/released new assets (e.g., code, data, models)…
+    1. (a)
+      Did you mention the license of the new assets (e.g., as part of your code submission)? Yes
+    2. (b)
+      Did you include the new assets either in the supplemental material or as a url (to, e.g., GitHub or Hugging Face)? Yes
+  6. 6.
+    If you used crowdsourcing or conducted research with human subjects…
+    1. (a)
+      Did you include the full text of instructions given to participants and screenshots, if applicable? NA
+    2. (b)
+      Did you describe any potential participant risks, with links to Institutional Review Board (irb) approvals, if applicable? NA
+    3. (c)
+      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
+  7. 7.
+    If you included theoretical results…
+    1. (a)
+      Did you state the full set of assumptions of all theoretical results? NA
+    2. (b)
+      Did you include complete proofs of all theoretical results? NA
+Appendix B Training across data fractions
+TangleNAS is more robust across dataset fractions for network weights and architecture optimization as seen from Tables 9, 10, 11, and 12.
+Search Type Optimizer Train portion CIFAR-10 (%) CIFAR-100 (%)
+Single-Stage TangleNAS 50% $97.715\pm 0.088$ $82.538\pm 0.118$
+80% $\mathbf{97.872\pm 0.054}$ $\mathbf{82.668\pm 0.161}$
+Two-Stage SPOS+RS 50% $97.680\pm 0.026$ $82.537\pm 0.280$
+80% $97.767\pm 0.024$ $82.210\pm 0.142$
+SPOS+ES 50% $97.77\pm 0.038$ $82.354\pm 0.120$
+80% $97.642\pm 0.023$ $82.517\pm 0.114$
+Table 9: Evaluation on the AutoFormer-T space for CIFAR-10 and CIFAR-100.
 Search Type Optimizer Train portion Supernet Accuracy (%)
-CIFAR10 CIFAR100 ImageNet16-120
-Single-Stage DrNAS 50% WS $\mathbf{94.36\pm 0.00}$ $72.245\pm 0.7315$ $\mathbf{46.37\pm 0.00}$
-80% $\mathbf{94.36\pm 0.00}$ $71.1525\pm 0.697$ $\mathbf{46.37\pm 0.00}$
+CIFAR-10 CIFAR-100 ImageNet16-120
+Single-Stage DrNAS 50% WS $\mathbf{94.36\pm 0.000}$ $72.245\pm 0.732$ $\mathbf{46.37\pm 0.00}$
+80% $\mathbf{94.36\pm 0.00}$ $71.153\pm 0.697$ $\mathbf{46.37\pm 0.00}$
 TangleNAS 50% WE $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$
 80% $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$
-Two-Stage SPOS+RS 50% WE $89.107\pm 0.88392$ $56.865\pm 2.5968$ $31.665\pm 1.1460$
-80% $87.778\pm 2.4462$ $53.68\pm 4.1736$ $30.5449\pm 3.6425$
-SPOS+ES 50% WE $87.133\pm 2.6050$ $56.463\pm 2.3417$ $29.7849\pm 3.0149$
-80% $89.095\pm 0.8250$ $56.363\pm 4.7244$ $30.935\pm 3.54648$
-Table 10: Comparison of test accuracies of single and two stage methods with WS and WE on NB201 search space
-Table 11: Evaluation on toy cell-based search space on fashion-mnist dataset
-Search Type Optimizer Train portion Supernet type Test acc
-Single-Stage DrNAS 50% WS $91.19\pm 0.0490$
-80% $91.125\pm 0.0334$
+Two-Stage SPOS+RS 50% WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.146$
+80% $87.778\pm 2.446$ $53.68\pm 4.174$ $30.545\pm 3.643$
+SPOS+ES 50% WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.015$
+80% $89.095\pm 0.825$ $56.363\pm 4.724$ $30.935\pm 3.546$
+Table 10: Comparison of test accuracies of single and two-stage methods with WS and WE on NB201 search space.
+Search Type Optimizer Train portion Supernet type Test acc (%)
+Single-Stage DrNAS 50% WS $91.19\pm 0.049$
+80% $91.125\pm 0.033$
 TangleNAS 50% WE $\mathbf{91.3\pm 0.023}$
-80% $91.065\pm 0.1633$
-Two-Stage SPOS+RS 50% WE $90.68\pm 0.25345$
-80% $90.687\pm 0.10969$
-SPOS+ES 50% WE $90.3175\pm 0.2233$
-80% $90.595\pm 0.2188$
+80% $91.065\pm 0.163$
+Two-Stage SPOS+RS 50% WE $90.68\pm 0.253$
+80% $90.687\pm 0.110$
+SPOS+ES 50% WE $90.318\pm 0.223$
+80% $90.595\pm 0.219$
 Optimum - - - 91.63
-Search Type Optimizer Train portion Supernet type Test acc
-Single-Stage DrNAS 50% WS 10%
-80% 10%
+Table 11: Evaluation on toy cell-based search space on Fashion-MNIST dataset.
+Search Type Optimizer Train portion Supernet type Test acc (%)
+Single-Stage DrNAS 50% WS 10
+80% 10
 TangleNAS 50% WE $\mathbf{83.020\pm 0.000}$
-80% 82.495±.0.46074plus-or-minus82.495.0.4607482.495$\pm$.0.4607482.495 ± .0.46074
-Two-Stage SPOS+RS 50% WE $81.253\pm 0.67174$
-80% $81.345\pm 0.38336$
-SPOS+ES 50% WE $81.890\pm 0.8002$
-80% $82.322\pm 0.60356$
+80%   $82.495\pm.0.461$
+Two-Stage SPOS+RS 50% WE $81.253\pm 0.672$
+80% $81.345\pm 0.383$
+SPOS+ES 50% WE $81.890\pm 0.800$
+80%    $82.322\pm 0.604$
 Optimum - - - 84.41
-Table 11: Evaluation on toy cell-based search space on fashion-mnist dataset
-Table 12: Evaluation on toy conv-macro search space on CIFAR10 dataset
-Search Type Optimizer Train portion Supernet CIFAR10 ImageNet
-Single-Stage TangleNAS 50% WE $\mathbf{2.556\pm 0.03406}$ $\mathbf{25.69}$
-80% $2.67\pm 0.0756$ 25.742
-DrNAS 50% WS $2.625\pm 0.0750$ 26.29
-80% $2.580\pm 0.02798$ $\mathbf{25.67}$
-Two-Stage SPOS+RS 50% WE $2.965\pm 0.07193$ 27.114
-80% $2.965\pm 0.07193$ 27.114
-SPOS+RE 50% WE $3.20000\pm 0.06557$ 27.424
-80% $3.00249\pm 0.037052$ 26.76
-Table 13: Comparison of test errors of single and two stage methods with WS and WE on DARTS search space
-Appendix B Search and Training Times
-B.1 NB201 and DARTS
-Search Type Optimizer Supernet Search Time (hrs)
-Single-Stage DrNAS WS 20.9166
-TangleNAS WE 20.3611
-Two-Stage SPOS+RS WE 29.51388
-SPOS+ES WE 26.74721
-Table 14: Comparison of search times NB201
-Search Type Optimizer Supernet GPU MemSearch Time(hrs)
-Single-Stage DrNAS WS 81.08% 29.4444
-TangleNAS WE 52.26% 25.472
-Two-Stage SPOS+RS WE 25.66% 18.7444
-SPOS+ES WE 25.66% 14.794
-Table 15: Comparison of test errors of single and two-stage methods on the DARTS search space
-Appendix C Search Space details
+Table 12: Evaluation on toy conv-macro search space on CIFAR-10 dataset.
+Search Type Optimizer Train portion Supernet CIFAR-10 (%) ImageNet (%)
+Single-Stage TangleNAS 50% WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.69}$
+80% $2.67\pm 0.076$ 25.742
+DrNAS 50% WS $2.625\pm 0.075$ 26.29
+80% $2.580\pm 0.028$ $\mathbf{25.67}$
+Two-Stage SPOS+RS 50% WE $2.965\pm 0.072$ 27.114
+80% $2.965\pm 0.072$ 27.114
+SPOS+ES 50% WE $3.200\pm 0.066$ 27.424
+80% $3.002\pm 0.037$ 26.76
+Table 13: Comparison of test errors of single and two-stage methods with WS and WE on DARTS search space.
+Appendix C Comparison with different gradient-based optimizers
+We also compare DrNAS against different gradient-based NAS optimizers in Tables 15, 14, and 16 on the toy cell-based, toy macro and the AutoFormer search spaces.
+We observe that DrNAS outperforms GDAS and DARTS on all of these search spaces, showing its robust nature.
+Optimizer Test-acc (%)
+TangleNAS+DrNAS 83.02
+SPOS+RS 81.25
+SPOS+ES 81.89
+TangleNAS+DARTS_v1 81.61
+TangleNAS+DARTS_v2 81.49
+TangleNAS+GDAS 10 (degenerate)
+Table 14: Comparison of DrNAS with other gradient-based optimizers on the toy conv-macro search space on CIFAR-10.
+Optimizer Test-acc (%)
+TangleNAS+DrNAS 90.930
+SPOS+RS 90.688
+SPOS+ES 90.595
+TangleNAS+DARTS_v1 89.905
+TangleNAS+DARTS_v2 90.747
+TangleNAS+GDAS 90.618
+Table 15: Comparison of DrNAS with other gradient-based optimizers on the toy cell-based search space on FashionMNIST dataset.
+Optimizer CIFAR-10 CIFAR-100
+TangleNAS+DrNAS 97.872 $\pm$ 0.054 82.668 $\pm$ 0.161
+SPOS+RS 97.767 $\pm$ 0.024 82.210 $\pm$ 0.142
+SPOS+ES 82.518 $\pm$ 0.114 82.518 $\pm$ 0.114
+TangleNAS+DARTS_v1 97.672 $\pm$ 0.040 82.107 $\pm$ 0.392
+TangleNAS+GDAS 97.45 $\pm$ 0.096 82.120 $\pm$ 0.281
+Table 16: Comparison of DrNAS with other gradient-based optimizers on the AutoFormer-T space for CIFAR-10 and CIFAR-100.
+Appendix D Search Space details
 Toy cell space
-This particular search space takes its inspiration from the prominently used Differentiable Architecture Search (DARTS) space, and is composed of diminutive triangular cells, with each edge offering four choices of operations: (a) Separable 3x3 Convolution, (b) Separable 5x5 Convolution, (c) Dilated 3x3 Convolution, and (d) Dilated 5x5 Convolution.
-The macro-architecture of the model comprises of three cells of types reduction, normal, and reduction again stacked together.
-Notably, we entangle the 3x3 and 5x5 kernel weights for each operation type, i.e., separable convolutions and dilated convolutions.
-We evaluate these search spaces and their architectures on the Fashion-MNIST dataset by creating a small benchmark which we release here.
+This particular search space takes its inspiration from the prominently used Differentiable Architecture Search (DARTS) space, and is composed of diminutive triangular cells, with each edge offering four choices of operations: (a) Separable $\times3$ Convolution, (b) Separable $\times5$ Convolution, (c) Dilated $\times3$ Convolution, and (d) Dilated $\times5$ Convolution.
+The macro-architecture of the model comprises three cells of the types reduction, normal, and reduction again stacked together.
+Notably, we entangle the $\times3$ and $\times5$ kernel weights for each operation type, i.e., separable convolutions and dilated convolutions.
+We evaluate these search spaces and their architectures on the Fashion-MNIST dataset by creating a small benchmark, which we release here.
 Toy conv-macro space
-This toy space draws its inspiration from MobileNet-like spaces where we search for the number of channels and the kernel size of convolutional layers in a network (also called macro search space) for four convolutional layers.
+This toy space draws its inspiration from MobileNet-like spaces where we search for the number of channels and the kernel size of convolutional layers in a network (also referred to as a macro search space) for four convolutional layers.
 Every convolutional layer has a choice of three kernel sizes and number of channels.
 See Table 17 for more details.
-We evaluate this search space on the CIFAR10 dataset by creating a small benchmark which we release here.
+We evaluate this search space on the CIFAR-10 dataset by creating a small benchmark, which we release here.
 Architectural Parameter Choices
@@
 Number of channels (layer 4) 64, 128, 256
-Table 16: Toy Convolutional-Macro Search Space
-NASBench201 Space
-The NASBench201 [Dong and Yang, 2020] has a single cell type which is stacked 5 times.
-Each cell has 4 nodes and each node are connected by operations from amongst skip connection, conv 3x3, conv 1x1, max-pool or average-pool.
-Opportunities for weight entanglement in this space is limited and we entangle the conv 3x3 and conv 1x1 kernel on every edge for every cell, thus obtaining parameter savings over traditional weight sharing.
+Table 17: Toy Convolutional-Macro Search Space.
+NAS-Bench-201 Space
+The NAS-Bench-201 search space (dong2020bench) has a single cell type which is stacked 15 times, with residual blocks (he-cvpr16a) after the fifth and tenth cells.
+Each cell has 4 nodes, with each node connected to the previous ones by operations chosen from skip connection, $\times3$ convolution, $\times1$ convolution, $\times3$ average-pooling, and zeroize, which zeros out the input feature map.
+There is limited scope for weight-entanglement in this space, given that only two of the five candidate operations contain learnable parameters.
+In this case, we entangle the $\times3$ and $\times1$ convolution kernels on every edge for every cell, thus obtaining parameter savings over traditional weight sharing.
+Architectural Parameter Choices
+Embedding Dimension 384, 576, 768
+Number of Heads 6, 8, 12
+MLP Expansion Ratio 2, 3, 4
+Number of Layers 5,6,7
+Table 18: Choices for Language Model Space.
 DARTS Search Space
-The DARTS [Liu et al., 2019] search space has a cell with 13 possible edges stacked 8 times.
-The nodes of the cell are connected by operations amongst :none, max_pool_3x3, avg_pool_3x3, skip_connect, sep_conv_5x5, sep_conv_3x3, dil_conv_5x5, dil_conv_3x3.
+The DARTS (liu-iclr19a) search space has two kinds of cells - normal and reduction.
+Reduction cells double the number of channels in its outputwhile halving the width and height of the feature maps.
+The supernet consists of 8 cells stacked sequentially, while the discretized network stacks 20 cells.
+Reduction cells are placed at 1/3 and 2/3 of the total depth of the network.
+Each cell takes two inputs - one each from the outputs of the two previous cells.
+There are 14 edges in each cell, with each edge in the supernet consisting of 8 candidate operations as follows: $\times3$ max pooling, $\times3$ average pooling, skip connect, $\times3$ separable convolutions, $\times5$ separable convolutions, $\times3$ dilated convolutions, $\times5$ dilated convolutions, and none (no operation).
+Architectural Parameter Choices
+Kernel sizes (all layers) 3, 5, 7
+Channel expansion (all layers) 3,4,6
+Number of blocks 2,3,4
+Table 19: MobileNetV3 Search Space.
 AutoFormer Space and Language Model Space
-We present the details of the AutoFormer Space and the Language Model space is Table 19 and Table 19 respectively.
-Table 17: Choices for AutoFormer-T Search Space
-Table 18: Choices for Language Model Space
+We present the details of the AutoFormer Space and the Language Model space is Tables 20 and 18, respectively.
 Architectural Parameter Choices
@@
 Number of heads (per layer) 3, 4
-Architectural Parameter Choices
-Embedding Dimension 384, 576, 768
-Number of Heads 6, 8, 12
-MLP Expansion Ratio 2, 3, 4
-Number of Layers 5,6,7
-Architectural Parameter Choices
-Kernel sizes (all layers) 3, 5, 7
-Channel expansion (all layers) 3,4,6
-Number of blocks 2,3,4
-Table 17: Choices for AutoFormer-T Search Space
-Table 18: Choices for Language Model Space
-Table 19: MobileNet Search Space
-Appendix D Details on Method
+Table 20: Choices for AutoFormer-T Search Space.
+Appendix E Methodological details
+Combi-Superposition
 Traditional cell-based search spaces primarily consider independent operations (e.g., convolution or skip).
-One-shot differentiable optimizers thus have their mixture operations tailored to these search spaces which are not general enough to be applied to macro-level architectural parameters.
+One-shot differentiable optimizers thus have their mixture operations tailored to these search spaces, which are not general enough to be applied to macro-level architectural parameters.
 Consider, e.g., the task of searching for the embedding dimension and the expansion ratio for a transformer.
-Here, a single operation, i.e., a linear expansion layer, has two different architectural parameters - one corresponding to the choice of embedding dimension and the other to the expansion ratio.
+Here, a single operation, i.e., a linear expansion layer, has two different architectural parameters – one corresponding to the choice of embedding dimension and the other to the expansion ratio.
 To adapt single-stage methods to these combined operation choices in the search space, we propose the combi-superposition operation.
-The combi-superposition operation simply takes the cross product of architectural parameters for the embedding dimension and expansion ratio and assigns its elements to every combination of these dimensions, allowing us to optimize jointly in this space without the need for a separate forward pass for each combination.
-Every combination maps to a unique sub-matrix of the operator weight matrix, indexed using both the embedding dimension and the expansion ratio.
+The combi-superposition operation simply takes the cross product of architectural parameters for the embedding dimension and expansion ratio and assigns its elements to every combination of these dimensions.
+This allows us to optimize jointly in this space without the need for a separate forward pass for each combination.
+Every combination maps to a unique sub-matrix of the operator weight matrix, indexed by both the embedding dimension and the expansion ratio.
 To address shape mismatches of the different operation weights during forward passes, every sub-matrix is zero-padded to match the shape of the largest matrix.
-See Algorithm 2 for more details, and Figure 2(b) for an overview of the idea.
+See Algorithm 2 for more details, and Figure 3 for an overview of the idea.
 Algorithm 2 Combi-Superposition Operation
+For ease of presentation, we show the algorithm for superimposing along two dimensions.
+However, in practice, we can superimpose an arbitrary number of dimensions (e.g., four dimensions in our experiments with NanoGPT).
 $embed\_dim\leftarrow[e_{1},e_{2},e_{3},\ldots,e_{n}]$ {Choices for embedding dimension}
@@
 $b_{mix}\leftarrow\textbf{0}$
-  $\alpha,\beta\leftarrow\text{normalize}(\alpha),\text{normalize}(\beta)$Normalize architecture parameters (e.g., using softmax)}
-  $\alpha\beta\leftarrow\alpha\times\beta$Cross product of $\alpha$ and $\beta$ for dimension $n\times m$}
-  for $i\leftarrow 1$ to $n\times m$ do
-     $W\_i=W[:(embed\_dim[i]\times expansion\_ratio[i]),:embed\_dim[i]]$
-     $b\_i=b[:embed\_dim[i]\times expansion\_ratio[i]]$
-     $W_{mix}=W_{mix}+\alpha\beta[i]\times\text{PAD}(W_{i})$
-     $b_{mix}=b_{mix}+\alpha\beta[i]\times\text{PAD}(b_{i})$
+for $i\leftarrow 1$ to $n$ do
+  for $j\leftarrow 1$ to $m$ do
+   $W\_{ij}=W[:(embed\_dim[i]\times expansion\_ratio[j]),:embed\_dim[i]]$
+   $b\_{ij}=b[:embed\_dim[i]\times expansion\_ratio[j]]$
+   $W_{\text{mix}}\leftarrow W_{\text{mix}}+\text{normalize}(\alpha[i])\cdot\text{normalize}(\beta[j])\times\text{PAD}(W_{ij})$
+   $b_{\text{mix}}\leftarrow b_{\text{mix}}+\text{normalize}(\alpha[i])\cdot\text{normalize}(\beta[j])\times\text{PAD}(b_{ij})$
+  end for
 end for
@@
 return $Y$
-For completeness and for comparison with TangleNAS (Algorithm 1), we present Algorithm 3 which describes a generic two-stage method on a macro search space with WE.
-Also, vanilla single-stage methods on cell-based WS spaces follow Algorithm 4.
+For completeness and to facilitate comparison with TangleNAS (Algorithm 1), we present Algorithm 3, which describes a generic two-stage method on a macro search space with weight entanglement (WE).
+Additionally, vanilla single-stage methods on cell-based weight sharing (WS) spaces follow Algorithm 4.
 Algorithm 3 Weight Entanglement (Two-Stage)
 1:Input: $M\leftarrow$ number of cells, $N\leftarrow$ number of operations    $\mathcal{O}\leftarrow[o_{1},o_{2},o_{3},...o_{N}]$   $\mathcal{W}_{\max}\leftarrow\cup_{i-1}^{N}w_{i}$    $\eta=$ learning rate of $\mathcal{W}_{\max_{\mathcal{O}}}$
-2:  $Cell_{j}\leftarrow DAG(\mathcal{O_{j},\mathcal{W}_{\max_{j}}})$ /* defined for j=1...M */
-3:  $Supernet\leftarrow\cup_{i}^{M}Cell_{i}$
+[…]
+[…]
 4:/* example of forward propagation on the cell */
@@
 11: $\mathcal{W}_{\max}[:i]=\mathcal{W}_{\max}[:i]-\eta\nabla_{\mathcal{W}_{\max}}[:i]\mathcal{L}_{train}(\mathcal{W}_{\max})$
-12:   /* Search */ 13:  $Supernet^{*}\leftarrow$ pre-trained supernet 14:  $selected\_arch\leftarrow$ Evolutionary-Search($Supernet^{*}$)
+[…]
 Algorithm 4 Weight Sharing (Single-Stage)
@@
 $\sum_{i=1}^{N}f(\alpha_{i})=1$
-2:  $Cell_{i}\leftarrow DAG(\mathcal{O_{j},\mathcal{W_{O_{j}}}})$ /* defined for i=1...M */
-3:  $Supernet\leftarrow\cup_{i}^{M}Cell_{i}\cup\mathcal{A}$
+[…]
+[…]
 4:/* example of forward propagation on the cell */
@@
 13: $selected\_arch\leftarrow$ $\arg\max(\mathcal{A})$
-Appendix E Experimental Setup
-E.1 Toy Search Spaces
+Compatibility issues between weight-entanglement and gradient-based methods
+We address the incompatibility between gradient-based NAS methods and weight-entangled (WE) spaces as follows (where n refers to the number of operation choices):
+  • Gradient-based NAS methods do not share or entangle weights among competing operations, which increases their GPU memory footprint.
+We tackle this issue by adopting weight-entanglement from two-stage methods, thereby reducing the parameter size of the supernet from $\mathcal{O}$}(n) to $\mathcal{O}$}(1).
+  • Gradient-based NAS methods compute a weighted combination of output of the operations.
+Even after entangling the operation weights, this approach leads to increased GPU memory usage because all intermediate competing activations/features need to be preserved in memory.
+Additionally, this method does not scale well with an increase in the number of operation choices, as GPU memory consumption during forward propagation scales as $\mathcal{O}$}(n).
+To address this, we propose weight superposition, which computes the weighted combination directly within the space of entangled weights.
+  • In vanilla gradient-based methods, a forward pass is computed independently for each operation choice, resulting in a time and memory complexity of $\mathcal{O}$}(n).
+In contrast, TangleNAS computes only a single forward pass using the superimposed weights, reducing the cost of the forward pass to $\mathcal{O}$}(1).
+Details on Figure 1
+In the overview Figure 1, we use rounded squares to represent the input and output feature maps of a convolution.
+The colored cubes represent the convolutional kernels, while the colored rectangles denotes non-convolutional architectural choices.
+These two differ primarily in how their weights are entangled.
+Consider Figure 1 (a), which illustrates the forward pass of an input feature map through the candidate operations on an edge of the supernet in a two-stage weight-entanglement method (such as OFA).
+In two-stage methods, random paths through the supernet are sampled and trained in the first stage.
+In this figure, thick lines indicate the paths sampled in a given step.
+The weights of the operation choices, depicted in the figure by colored cubes and rectangles, overlap with one another, showing that all the operations use slices of a common, large weight matrix.
+We show three choices of convolutions, each of a different kernel size (say, $\times1$, $\times3$, and $\times5$), represented by cubes of varying sizes.
+The $\times1$ and $\times3$ convolutions use slices of the larger $\times5$ convolution as their weights, represented in the figure by nesting the smaller kernels inside the larger ones.
+Since only one operation is sampled at a time (in this case, the orange kernel), there is only one output feature map.
+This feature map, after global average pooling, is then passed through one of the three non-convolutional operation choices.
+The operation choices here may be the embedding dimension, for example, and different slices of the largest feedforward network are used for the choices of embedding dimensions.
+At the end of the first stage in Figure 1 (a), the supernet has been trained along different paths.
+In the second stage, paths are sampled from the trained supernet using black-box methods (such as random or evolutionary search) and evaluated on the dataset to obtain the optimal architecture, which we represent with optim arch.
+Now, consider Figure 1 (b), which represents the forward pass in a single-stage weight-sharing method (such as DARTS).
+As you can see, there is no nesting of the weights of the three convolutions or feedforward networks in this case, indicating that each has its own distinct set of weights.
+Naturally, this will incur more GPU memory usage, as more weights need to be stored and their gradients computed.
+The outputs from each of these convolutional (or feedforward) operations are then weighted by $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ (or $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$), which represent the architectural parameters of the operations.
+These weighted feature maps are then summed up to produce the output feature maps.
+The optimization loop of single-stage weight-sharing methods, shown in Figure 1(b), aims to find the optimal values for the architectural parameters $\alpha$ and $\beta$.
+We obtain the discretized model by selecting only the operations in the supernet with the highest values of these parameters at the end of this loop.
+Specifically, the convolutional kernel and non-convolutional operator (in this case, the feedforward network) with the maximum values of $\alpha$ and $\beta$ are represented as argmax($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$) and argmax($\beta_{1}$, $\beta_{2}$, $\beta_{3}$), respectively.
+The resulting discretized model, or optimal architecture, is denoted as optimal arch.
+Note that these methods do not require black-box search, as the optimal architecture can be obtained directly from the learned architectural parameters.
+In Figure 1 (c), we depict our hybrid framework, which utilizes both entangled (nested) weights and architectural parameters.
+This approach reduces GPU memory usage due to weight entanglement and allows for the optimal architecture to be obtained directly from the learned parameters, eliminating the need for a black-box search.
+Appendix F Experimental Setup
+F.1 Toy Search Spaces
+Below are the hyperparameter settings for the two toy spaces for TangleNAS and SPOS. All experiments were run on a single RTX-2080 GPU.
 Search Space Cell Space Conv Macro
@@
 Table 22: Configurations used in the SPOS experiments on Toy Spaces.
-E.2 Language Model
-We use the AdamW optimizer in all experiments pertaining to language modelling. Other hyperparameter choices are as in Table 22.
+F.2 Language Model
+We use the AdamW optimizer in all experiments related to language modeling.
+Other hyperparameter choices are as specified in Table 23.
+Below are the hyperparameter settings for TangleNAS.
+We run experiments on TinyStories on 8 RTX-2080 GPUs and experiments on OpenWebText on 8 A6000 GPUs.
 Search Space Small-LM
@@
 Weight decay 1e-1
-Table 22: Configurations used in DrNAS on the Language Model Spaces
-E.3 AutoFormer and OFA
-We use a 50%-50% train and validation split for the CIFAR10 and CIFAR100 dataset for the cell-based spaces and a 80%-20% for the weight-entangled spaces We use the official source code of AutoFormer available at code for all the AutoFormer experiments on CIFAR10 and CIFAR100.
-We closely follow the AutoFormer training pipeline and search space design.
-AutoFormer searched on three transformer sizes Autoformer-T (tiny), AutoFormer-S (small) and Autoformer-B (base).
-We currently restrict ourselves to Autoformer-T.
-For baselines like OFA and AutoFormer we follow their respective recipes to obtain the train-valid split for ImageNet-1k.
+Table 23: Configurations used in DrNAS on the Language Model Spaces.
+F.3 AutoFormer and OFA
+We use a 50%-50% train and validation split for the CIFAR-10 and CIFAR-100 datasets for the cell-based spaces and a 80%-20% for the weight-entangled spaces.
+We use the official source code of AutoFormer available at code autoformer for all the AutoFormer experiments on CIFAR-10 and CIFAR-100, closely following the AutoFormer training pipeline and search space design.
+AutoFormer explored three transformer sizes: Autoformer-T (tiny), AutoFormer-S (small), and Autoformer-B (base).
+We restrict ourselves to Autoformer-T.
+For baselines like OFA and AutoFormer, we follow their respective recipes to obtain the train-validation split for ImageNet-1k.
 Our models were trained on 2xA100s with the same effective batch-size as AutoFormer.
-For MobileNetV3 from once-for-all we use the same training hyperparameters as the baseline as here (in addition to architectural parameters same as Table 23) .
-E.3.1 AutoFormer Fine-tuning
-C10 and C100 pretrained supernet
-We fine-tune the C10 and C100 selected networks (after inheriting them from the supernet) for 500 epochs.
-We set the learning rate to 1e-3, the warmup epochs to 5 , the warmup learning rate to 1e-6 and the min learning rate to 1e-5.
-All other hyperparameters are set same as Section E.3.
+For MobileNetV3 from Once-for-All, we use the same training hyperparameters as the baseline found here (in addition to architectural parameters same as Table 24).
+We run experiments on CIFAR-10, CIFAR-100, and ImageNet1-k on 4 A100 GPUs.
+F.3.1 AutoFormer Fine-tuning
+CIFAR-10 and CIFAR-100 pretrained supernet
+We fine-tune the CIFAR-10 and CIFAR-100 selected networks (after inheriting them from the supernet) for 500 epochs.
+We set the learning rate to 1e-3, the warmup epochs to 5, the warmup learning rate to 1e-6, and the minimum learning rate to 1e-5.
+All other hyperparameters are set the same as in Appendix F.3.
 ImageNet pretrained supernet
-We follow the DeiT [Touvron et al., 2021] finetuning pipeline as used in AutoFormer.
-Mainly we set the epochs to 1000, the warmup-epochs to 5 , the scheduler to cosine, the mixup to 0.8, the smoothing to 0.1, the weight-decay to 1e-4, the batch-size 64, the optimizer to sgd, the learning rate to 0.01 and the warmup-lr to 0.0001 for all datasets.
-E.4 NB201 and DARTS
+We follow the DeiT (touvron-icml21) fine-tuning pipeline as used in AutoFormer, to finetune on downstream tasks.
+Specifically, we set the epochs to 1000, the warmup epochs to 5, the scheduler to cosine, the mixup to 0.8, the smoothing to 0.1, the weight decay to 1e-4, the batch size to 64, the optimizer to SGD, the learning rate to 0.01 and the warmup learning rate to 0.0001 for all datasets.
+Fine-tuning is performed on 8 RTX-2080 GPUs.
+F.4 NB201 and DARTS
 For single-stage optimizers, the supernet was trained with four different seeds.
-The supernet with the best validation performance from these four was discretized to obtain the final model, which is then trained from scratch four times to obtain the results shown in the table.
-For two-stage methods, yet again, we train the supernet four times, and perform Random Search (RS) and Evolutionary Search (ES) on each of them.
-The best model obtained across all four supernets for both methods are then trained from scratch with four seeds to compute the final results.
-For each DrNAS we follow the same recipe as suggested by the authors across all search spaces.
-To accommodate for multiple training recipes, we write a configurable training pipeline.
-The configurations for DrNAS, and SPOS are shown in Table 23 and Table 24, respectively.
+The supernet with the best validation performance among these four was discretized to obtain the final model, which was then trained from scratch four times to obtain the results shown in the table.
+For two-stage methods, we again train the supernet four times and perform Random Search (RS) and Evolutionary Search (ES) on each one.
+The best model obtained across all four supernets for both methods was then trained from scratch with four seeds to compute the final results.
+For DrNAS, we follow the same procedure as suggested by the authors across all search spaces.
+To accommodate multiple training recipes, we have developed a configurable training pipeline.
+The configurations for DrNAS and SPOS are shown in Tables 24 and 25, respectively.
+We run our search on a single RTX-2080 for both DARTS and NB201, and train the DARTS architectures from scratch on 8 RTX-2080 GPUs.
 Search Space DARTS NB201
@@
 Table 25: Configurations used in the SPOS experiments.
-Appendix F Optimal architectures derived
-F.1 DARTS
+Appendix G Optimal architectures derived
+G.1 DARTS
 Figure 7: DRNAS with WE normal cell
 Figure 8: DRNAS with WE reduction cell
-F.2 Small LM
+G.2 Small LM
 num_layers: 7, embed_dim: 768, num_heads: [12, 12, 12, 12, 12, 12, 12], mlp_ratio: [3, 4, 4, 4, 4, 4, 4]
-F.3 AutoFormer
-F.3.1 CIFAR10
+G.3 AutoFormer
+G.3.1 CIFAR-10
 50%50% split
-mlp_ratio:$[4,4,4,4,4,4,4,4,4,3.5,4,4,3.5,4]$, num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4]44444444444444[4,4,4,4,4,4,4,4,4,4,4,4,4,4][ 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 ], num_layers: 14 embed_dim: 216
+mlp_ratio:$[4,4,4,4,4,4,4,4,4,3.5,4,4,3.5,4]$, num_heads:$[4,4,4,4,4,4,4,4,4,4,4,4,4,4]$, num_layers: 14 embed_dim: 216
 80%-20% split
-mlp_ratio:$[4,4,4,4,4,4,4,4,3.5,4,4,4,3.5,3.5]$, num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4]44444444444444[4,4,4,4,4,4,4,4,4,4,4,4,4,4][ 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 , 4 ], depth: 14, embed_dim: 240
-F.3.2 CIFAR100
+mlp_ratio:$[4,4,4,4,4,4,4,4,3.5,4,4,4,3.5,3.5]$, num_heads:$[4,4,4,4,4,4,4,4,4,4,4,4,4,4]$, depth: 14, embed_dim: 240
+G.3.2 CIFAR-100
 50%-50% split
@@
 mlp_ratio:[3.5,4,4,4,4,4,4,4,4,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4, depth: 14, embed_dim: 240
-F.3.3 ImageNet1-k
+G.3.3 ImageNet1-k
 mlp_ratio:[4,4,4,4,4,4,4,4,4,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4, depth: 14, embed_dim: 240
-F.4 MobileNetV3
+G.4 MobileNetV3
 Kernel_sizes:[7,5,5,7,5,5,7,7,5,7,7,7,5,7,7,5,5,7,7,5],Channel_expansion_factor:[6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6], Depths : [4, 4, 4, 4, 4]
-F.5 Toy Spaces
-F.5.1 Toy cell (our best architecture)
+G.5 Toy Spaces
+G.5.1 Toy cell (our best architecture)
 50%-50% split
@@
 Genotype(normal=[(’dil_conv_3x3’, 0), (’dil_conv_5x5’, 0), (’sep_conv_3x3’, 1)], normal_concat=range(1, 3), reduce=[(’sep_conv_3x3’, 0), (’sep_conv_5x5’, 0), (’dil_conv_3x3’, 1)], reduce_concat=range(1, 3))
-F.6 Toy conv-macro (our best architecture)
+G.6 Toy conv-macro (our best architecture)
 50%-50% split:
@@
 Channels = [32, 64, 128, 64], Kernel Sizes = [5, 5, 7, 7].
-Appendix G Architecture Representation Analysis
+Appendix H Architecture Representation Analysis
 CKA
-Centered Kernel Alignment (CKA) [Kornblith et al., 2019] is a metric, based on the Hilbert- Schmidt Independence Criterion (HSIC).
-This metric is primarily designed to model similarity between representations in neural networks.
-In this section we study the CKA between structurally identical layers in the inherited, fine-tuned and retrained networks in the AutoFormer-T space.
-Precisely, the goal here is to study how similar the inherited and fine-tuned representations are to the actual model trained from scratch.
-Table 25 Presents the average CKA values for a fixed subsample of the CIFAR10 and CIFAR100 dataset.
-We observe that the representations learned by the single-stage supernet are more similar to the architecture finetuned or trained from scratch.
-This further emphasizes the potential issues with using the inherited weights from the supernet as proxy for search for two-stage methods as highlighted in [Xu et al., 2022].
+Centered Kernel Alignment (CKA) (kornblith2019similarity) is a metric based on the Hilbert-Schmidt Independence Criterion (HSIC).
+It is designed to model the similarity between representations in neural networks.
+In this section, we analyze the CKA between structurally identical layers in the inherited, fine-tuned, and retrained networks in the AutoFormer-T space.
+Specifically, the aim is to assess how similar the inherited and fine-tuned representations are to those of the models trained from scratch.
+Table 26 presents the average CKA values for a fixed subset of the CIFAR-10 and CIFAR-100 datasets.
+We find that the representations learned by the single-stage supernet are more similar to those of the architectures that are fine-tuned or trained from scratch.
+This observation underscores potential issues with using inherited weights from the supernet as a proxy for search in two-stage methods, as noted by xu2022analyzing.
 Model Inherit v/s Retrain Fine-Tune v/s Retrain
-CIFAR10 CIFAR100 CIFAR10 CIFAR100
-TangleNAS 0.4630 0.5853 0.57125 0.65275
-SPOS+ES 0.45812 0.57932 0.569374 0.6309
-SPOS+RS 0.44124 0.583 0.5797 0.638948
-Table 25: CKA correlation between layers
-Architecture Search-Type Loss Perplexity Params Inference Time
-GPT-2 Manual 3.0772 21.69 123.59M 113.3s
-TangleNAS Automated 2.9038 18.243 116.51M 102.5s
-  1. 1.
-    Best Practices for Releasing Code
-    For all experiments you report:
-    1. (a)
-      Did you release code for the training pipeline used to evaluate the final architectures? Yes
-    2. (b)
-      Did you release code for the search space? Yes
-    3. (c)
-      Did you release the hyperparameters used for the final evaluation pipeline, as well as random seeds? Yes
-    4. (d)
-      Did you release code for your NAS method? Yes
-    5. (e)
-      Did you release hyperparameters for your NAS method, as well as random seeds? Yes
-  2. 2.
-    Best practices for comparing NAS methods
-    1. (a)
-      For all NAS methods you compare, did you use exactly the same NAS benchmark, including the same dataset (with the same training-test split), search space and code for training the architectures and hyperparameters for that code?
-Yes
-    2. (b)
-      Did you control for confounding factors (different hardware, versions of DL libraries, different runtimes for the different methods)? Yes
-    3. (c)
-      Did you run ablation studies? Yes
-    4. (d)
-      Did you use the same evaluation protocol for the methods being compared? Yes
-    5. (e)
-      Did you compare performance over time? Yes
-    6. (f)
-      Did you compare to random search? Yes. Black box random search is very expensive in our case, hence we use the supernet from single-path-one-shot(SPOS) for random search.
-    7. (g)
-      Did you perform multiple runs of your experiments and report seeds? Yes
-    8. (h)
-      Did you use tabular or surrogate benchmarks for in-depth evaluations? Yes
-  3. 3.
-    Best practices for reporting important details
-    1. (a)
-      Did you report how you tuned hyperparameters, and what time and resources this required? N/A. We did not tune any hyperparameters.
-    2. (b)
-      Did you report the time for the entire end-to-end NAS method (rather than, e.g., only for the search phase)? Yes
-    3. (c)
-      Did you report all the details of your experimental setup? Yes
-Eer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">
+CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
+TangleNAS 0.4630 0.5853 0.5712 0.6527
+SPOS+ES 0.4581 0.5793 0.5694 0.6309
+SPOS+RS 0.4412 0.583 0.5797 0.6389
+Table 26: CKA correlation between layers.
+Appendix I Limitations and Future Work
+Currently, TangleNAS is designed to optimize a single objective, such as a chosen performance metric.
+However, in practice, there may be multiple objectives of interest, including hardware efficiency, robustness, and fairness (dooley-neurips23).
+Additionally, it would be valuable to explore and apply our findings across various applications in computer vision and natural language processing.